### v1

# Function theory on the annulus in the dp-norm

Abstract.
In this paper we shall use realization theory, a favourite technique of Rien Kaashoek, to prove new results about a class of holomorphic functions on an annulus
$R_{\delta}\stackrel{{\scriptstyle\rm def}}{{=}}\{z\in\mathbb{C}:\delta<|z|<1\},$
where $10<\delta<$ 1. The class of functions in question arises in the early work of R. G. Douglas and V. I. Paulsen on the rational dilation of a Hilbert space operator $T$ to a normal operator with spectrum in $\partial R_{\delta}$. Their work suggested the following norm $\|\cdot\|_{\mathrm{dp}}$ on the space $\mathrm{Hol}(R_{\delta})$ of holomorphic functions on $R_{\delta}$,
$\|\varphi\|_{\mathrm{dp}}\stackrel{{\scriptstyle\rm def}}{{=}}\sup\{\|\varphi(T)\|:\|T\|\leq 1,\|T^{-1}\|\leq 1/\delta\ \text{and}\ \sigma(T)\subseteq R_{\delta}\}.$
By analogy with the classical Schur class of holomorphic functions $\mathcal{S}$ with supremum norm at most $1$ on the disc $\mathbb{D}$, it is natural to consider the dp-Schur class $\mathcal{S}_{\mathrm{dp}}$ of holomorphic functions of dp-norm at most $1$ on $R_{\delta}$.
Our central result is a generalization of the classical realization formula, for $\varphi\in\mathcal{S}$,
$\varphi(z)=A+Bz(1-Dz)^{-1}C,\ \text{for}\ z\in\mathbb{D},$
to functions from $\mathcal{S}_{\mathrm{dp}}$. A second result is a Pick interpolation theorem for functions in $\mathcal{S}_{\mathrm{dp}}$ that is analogous to Abrahamse’s Interpolation Theorem for bounded holomorphic functions on a multiply-connected domain. For a tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of distinct interpolation nodes in $R_{\delta}$, we introduce a special set $\mathcal{G}_{\mathrm{dp}}(\lambda)$ of positive definite $n\times n$ matrices, which we call DP Szegő kernels. The DP Pick problem $\lambda_{j}\mapsto z_{j},j=1,\dots,n$, is shown to be solvable if and only if,
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0\;\text{ for all}\;g\in\mathcal{G}_{\mathrm{dp}}(\lambda).$
We prove further that a solvable DP Pick problem has a solution which is a rational function with a finite-dimensional model, an intriguing result which opens up the possibility of a theory of extremal functions from $\mathcal{S}_{\mathrm{dp}}$ analogous to the theory of finite Blaschke products.
Key words and phrases:
Holomorphic functions; Hilbert space model; annulus; Crouzeix conjecture; Douglas-Paulsen family
2020 Mathematics Subject Classification:
47B99, 30E05, 32A26
Partially supported by National Science Foundation Grants DMS 1361720 and 1665260, the Engineering and Physical Sciences Research Council grant EP/N03242X/1.
1. Introduction
It is our honour to contribute to this memorial issue for Marinus Kaashoek, who was a prolific and influential operator theorist throughout a long career. A constant thread in his research over several decades was the power of realization theory applied to a wide variety of problems in analysis. Among his many contributions in this area we mention his monograph [8], written with his longstanding collaborators Israel Gohberg and Harm Bart, which was an early and influential work in the area, and his more recent papers and book, including [15, 14, 16]. Realization theory uses explicit formulae for functions in terms of operators on Hilbert space to prove function-theoretic results. In this paper we continue along the Bart-Gohberg-Kaashoek path by exploiting realization theory to prove new results about a class of holomorphic functions which was first encountered by R. G. Douglas and V. I. Paulsen in a study of rational dilation on the annulus.
For any open set $\Omega$ in the plane, ${\rm Hol}(\Omega)$ will denote the set of holomorphic functions on $\Omega$ and ${\rm H}^{\infty}(\Omega)$ will denote the Banach algebra of bounded holomorphic functions on $\Omega$, equipped with the supremum norm $\|\varphi\|_{{\rm H}^{\infty}(\Omega)}=\sup_{z\in\Omega}|\varphi(z)|$. Let $\mathcal{S}(\Omega)$ denote the class $\{\varphi\in{\rm H}^{\infty}(\Omega):\|\varphi\|_{{\rm H}^{\infty}(\Omega)}\leq 1\}$. The classical Schur class, $\mathcal{S}$, is the set $\mathcal{S}(\mathbb{D})$.
We recall the extensively-studied Pick interpolation theorem [20] for bounded holomorphic functions on the open unit disc $\mathbb{D}$.
Theorem 1.1.
Let $\lambda_{1},\dots,\lambda_{n}\in\mathbb{D}$ be distinct and let $z_{1},\dots,z_{n}\in\mathbb{C}$. There exists $\varphi\in\mathcal{S}$ such that
$\varphi(\lambda_{j})=z_{j}\qquad\text{for}\ j=1,\ldots,n,$
if and only if,
$\begin{bmatrix}\displaystyle\frac{1-\overline{z_{i}}z_{j}}{1-\overline{\lambda_{i}}\lambda_{j}}\end{bmatrix}_{i,j=1}^{n}\geq 0.$
Pick interpolation problems, with the unit disc replaced by other domains $\Omega$ in the plane, have also been much studied. In the event that $\Omega$ is a simply connected proper open subset of $\mathbb{C}$, with the aid of the conformal map $F:\Omega\to\mathbb{D}$, we can convert this problem into a classical Pick problem on $\mathbb{D}$ with interpolation data $F(\lambda_{j})\mapsto w_{j}$ for $j=1,\ldots,n,$ and then Pick’s theorem gives a criterion for the existence of $\varphi$ in terms of the positivity of the appropriate “Pick matrix”, which here is
$\begin{bmatrix}\displaystyle\frac{1-\overline{w_{i}}w_{j}}{1-\overline{F(\lambda_{i})}F(\lambda_{j})}\end{bmatrix}_{i,j=1}^{n}\geq 0.$
More generally, the Pick problem on a multiply connected domain was studied in the 1940s by Garabedian [17] and Heins [19]. Later, Sarason [21] and Abrahamse [1] treated the problem in terms of reproducing kernels, an approach that we follow in this paper. Abrahamse’s Theorem gives a solution to the Pick interpolation problem on any bounded domain $\Omega$ in the plane whose boundary consists of finitely many disjoint analytic Jordan curves. He showed that a Pick problem on $\Omega$ can be solved if and only if an infinite collection of Pick matrices are positive semi-definite. In the case of the annulus $R_{\delta}=\{z\in\mathbb{C}:\delta<|z|<1\}$, for a tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of distinct interpolation nodes in $R_{\delta}$, Abrahamse [1] described a family $\mathcal{G}(\lambda)$ of positive definite $n\times n$ matrices for which the following statement is true:
Theorem 1.2.
Let $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$ be distinct and let $z_{1},\dots,z_{n}\in\mathbb{C}$. There exists $\varphi\in\mathcal{S}(R_{\delta})$ such that
$\varphi(\lambda_{j})=z_{j}\qquad\text{for}\ j=1,\ldots,n,$
if and only if, for each $g\in\mathcal{G}(\lambda)$,
$\begin{bmatrix}(1-\overline{z_{i}}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}\geq 0.$
An alternative explicit choice of $\mathcal{G}(\lambda)$ for which Theorem 1.2 is true is described in [21, 2] as follows
$\mathcal{G}(\lambda)=\{\begin{bmatrix}g_{\rho}(\lambda_{i},\lambda_{j})\end{bmatrix}_{i,j=1}^{n}:\rho>0\},$
where
$g_{\rho}(\lambda_{i},\lambda_{j})=\sum_{m=-\infty}^{\infty}\frac{(\bar{\lambda}_{i}\lambda_{j})^{m}}{\rho+\delta^{2m}},\quad\text{for}\ 1\leq i,j\leq n.$
Another natural variant of Pick’s problem arises if one replaces the supremum norm on ${\rm Hol}(\Omega)$ by a different norm. For example, consider the Dirichlet space $\mathcal{D}$ of holomorphic functions $f$ on $\mathbb{D}$ such that $f^{\prime}$ is square integrable with respect to area measure on $\mathbb{D}$, with pointwise operations and the norm
[…]
where $m$ denotes area measure on the disc. The Dirichlet space is a Hilbert function space on $\mathbb{D}$ with reproducing kernel
$k_{\mathcal{D}}(\lambda,\mu)=-\frac{1}{\bar{\mu}\lambda}\log(1-\bar{\mu}\lambda).$
The Pick-type interpolation problem appropriate to this Hilbert function space is expressed in terms of its multiplier space $\mathcal{M}(\mathcal{D})$, which is defined to be the space of functions $\varphi$ on $\mathbb{D}$ such that $\varphi f\in\mathcal{D}$ for every $f\in\mathcal{D}$, with pointwise operations and the multiplier norm
$\|\varphi\|_{\mathcal{M}(\mathcal{D})}=\sup\{\|\varphi f\|_{\mathcal{D}}:f\in\mathcal{D},\|f\|_{\mathcal{D}}\leq 1\}.$
In this setting the corresponding Pick interpolation theorem is the following [3, Corollary 7.41]:
Theorem 1.3.
Let $\lambda_{1},\dots,\lambda_{n}\in\mathbb{D}$ be distinct and let $z_{1},\dots,z_{n}\in\mathbb{C}$. There exists $\varphi\in\mathcal{M}(\mathcal{D})$ such that $\|\varphi\|_{\mathcal{M}(\mathcal{D})}\leq 1$ and
$\varphi(\lambda_{j})=z_{j}\qquad\text{for}\ j=1,\ldots,n,$
if and only if
$\begin{bmatrix}(1-z_{i}\overline{z_{j}})k_{\mathcal{D}}(\lambda_{i},\lambda_{j})\end{bmatrix}_{i,j=1}^{n}\geq 0.$
An account of Pick theorems in the context of sundry different Hilbert function spaces, including $\mathcal{D},$ may be found in the book [3].
[…]
$\|\varphi\|_{\mathrm{dp}}=\sup_{X\in\mathcal{F}_{\mathrm{dp}}(\delta)}\|\varphi(X)\|,$ (1.4)
defined for $\varphi\in{\rm Hol}(R_{\delta})$. There is no guarantee that the quantity defined by equation (1.4) is finite. Accordingly, we introduce the associated Banach algebra
${\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})=\{\varphi\in{\rm Hol}(R_{\delta})\,:\,\|\varphi\|_{\mathrm{dp}}<\infty\}.$
In addition, we introduce the dp-Schur class, $\mathcal{S}_{\mathrm{dp}}$ [^2] In the notations $\|\cdot\|_{\mathrm{dp}}$ and $\mathcal{S}_{\mathrm{dp}}$ we suppress dependence on the parameter $\delta$., which is the set of functions $\varphi\in{\rm Hol}(R_{\delta})$ such that $\|\varphi\|_{\mathrm{dp}}\leq 1.$
An important step in the Douglas-Paulsen theory was the following estimate. If $X$ is a Douglas-Paulsen operator with parameter $\delta$, $\sigma(X)\subseteq R_{\delta}$ and $\varphi$ is a bounded holomorphic matrix-valued function on $R_{\delta}$ then
$\|\varphi(X)\|\leq\left(2+\frac{1+\delta}{1-\delta}\right)\sup_{z\in R_{\delta}}\|\varphi(z)\|.$ (1.5)
Hence, we see from equations (1.4) and (1.5) that $\|\varphi\|_{\mathrm{dp}}\leq\left(2+\frac{1+\delta}{1-\delta}\right)\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}$ for $\varphi\in{\rm Hol}(R_{\delta})$. On the other hand, trivially, $\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}\leq\|\varphi\|_{\mathrm{dp}}$, and so the dp and supremum norms on ${\rm Hol}(R_{\delta})$ are equivalent. Thus,
$\mathrm{H}^{\infty}(R_{\delta})=\mathrm{H}_{\mathrm{dp}}^{\infty}(R_{\delta})$
as sets. However, the reader should be aware that
$\|\cdot\|_{\mathrm{dp}}\neq\|\cdot\|_{{\rm H}^{\infty}(R_{\delta})}\ \text{andtherefore}\ \mathcal{S}_{\mathrm{dp}}\neq\mathcal{S}(R_{\delta}),$
a fact that Example 2.9 below demonstrates.
The power of inequality (1.5) is that it holds for all matrix-valued functions $\varphi$, a fact which allowed Douglas and Paulsen to show that if $T\in\mathcal{B}(\mathcal{H})$ is a Douglas-Paulsen operator, then there exists an invertible $S\in\mathcal{B}(\mathcal{H})$ such that
$\|S\|\|S^{-1}\|\leq 2+\frac{1+\delta}{1-\delta}$ (1.6)
and $STS^{-1}$ dilates to a normal operator with spectrum contained in the boundary $\partial R_{\delta}$. This result is a kind of Nagy dilation theorem for the annulus. In the scalar case a slightly stronger result than the inequality (1.5) had been obtained earlier by A. Shields [23, Proposition 23], with the smaller constant $2+\sqrt{\frac{1+\delta}{1-\delta}}$ on the right hand side. Shields asked whether the constant $2+\sqrt{\frac{1+\delta}{1-\delta}}$ could be replaced by a quantity that remains bounded as $\delta\to 1$. This question was answered in the affirmative by C. Badea, B. Beckermann and M. Crouzeix [7] and subsequently a better constant was established by M. Crouzeix [11].
Corresponding to the dp-Schur class there is a natural variant of the classical Pick interpolation problem, which we call the DP Pick problem: given $n$ distinct points $\lambda_{1},\ldots,\lambda_{n}$ in $R_{\delta}$ and $z_{1},\ldots,z_{n}\in\mathbb{C}$, does there exist a function $\varphi\in{\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ with $\|\varphi\|_{\mathrm{dp}}\leq 1$ such that
$\varphi(\lambda_{j})=z_{j}\qquad\text{for}\ j=1,\ldots,n?$ (1.7)
We shall show that there is a solvability criterion for this problem which is parallel to Abrahamse’s Theorem, but with $\mathcal{G}(\lambda)$ replaced by a collection $\mathcal{G}_{\mathrm{dp}}(\lambda)$ of kernels, which we now define.
Definition 1.8.
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$ and let $\lambda=(\lambda_{1},\dots,\lambda_{n})$. A DP Szegő kernel for the $n$-tuple $\lambda$ is a positive definite $n\times n$ matrix $g=[g_{ij}]$ such that
$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]\geq 0\ \ \text{ and }\ \ \left[\left(1-\overline{\frac{\delta}{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]\geq 0.$ (1.9)
The set of all DP Szegő kernels for the $n$-tuple $\lambda$ will be denoted by $\mathcal{G}_{\mathrm{dp}}(\lambda)$.
We observe that $\mathcal{G}_{\mathrm{dp}}(\lambda)$ consists of the gramians $[\langle e_{j},e_{i}\rangle]_{i,j=1}^{n}$ for all bases $e_{1},\dots,e_{n}$ of an $n$-dimensional Hilbert space $\mathcal{H}$ such that the operator $T$ on $\mathcal{H}$ defined by $Te_{j}=\lambda_{j}e_{j}$ for $j=1,\dots,n$ is a Douglas-Paulsen operator. This and related facts are described in Section 4.
The Pick interpolation theorem for the dp-norm on ${\rm Hol}(R_{\delta})$ is the following statement (which is Theorem 5.2 from the body of the paper).
Theorem 1.10.
Let $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$ be distinct and let $z_{1},\dots,z_{n}\in\mathbb{C}$. There exists $\varphi\in\mathcal{S}_{\mathrm{dp}}$ such that
$\varphi(\lambda_{j})=z_{j}\qquad\text{for}\ j=1,\ldots,n,$
if and only if, for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$ (1.11)
In Section 2 we compare the dp norm and the sup norm of a function in ${\rm Hol}(R_{\delta})$ and we point out a connection to the Crouzeix conjecture. In Section 3 we review the theory of models and realizations of holomorphic functions on $R_{\delta}$ with dp-norm at most $1$, see Theorem 3.9. In Section 4 we introduce DP-Szegő kernels on an $n$-tuple of points in $R_{\delta}$ and elaborate their relation to the Douglas-Paulsen class. In Section 5 we recall another approach to the solution of DP Pick problems given in [6, Theorem 9.46], and we show that solvable DP Pick problems have rational solutions. In Section 6 we consider an extremally solvable DP Pick problem $\lambda_{j}\mapsto z_{j}$ for $j=1,\dots,n$, and show that, for such a problem there is a rational solution $\varphi\in\mathcal{S}_{\mathrm{dp}}$ and there exists a Douglas-Paulsen operator $T$ with parameter $\delta$ which acts on an $n$-dimensional Hilbert space with $\sigma(T)=\{\lambda_{1},\dots,\lambda_{n}\}$ such that $\|\varphi\|_{\rm dp}=\|\varphi(T)\|=1,$ see Theorem 6.13.
2. The dp and sup norms on ${\rm Hol}(\mathbb{D})$ and ${\rm Hol}(R_{\delta})$
In this section we describe connections between the Banach algebra ${\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ and the Crouzeix conjecture. We will show in Proposition 2.11 that there is a large class of functions $\varphi\in{\rm Hol}(R_{\delta})$, such that
$\|\varphi\|_{\mathrm{dp}}=\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}.$
A simple example of a function $f\in{\rm Hol}(R_{\delta})$ for which the last relation fails is $f(z)=z+\frac{\delta}{z}$, which satisfies
$\|f\|_{\mathrm{dp}}=2\ \text{and}\ \|f\|_{{\rm H}^{\infty}(R_{\delta})}=1+\delta.$
By an elliptical domain we shall mean the domain in the complex plane bounded by an ellipse. As a standard elliptical domain we take the set
$G_{\delta}\stackrel{{\scriptstyle\rm def}}{{=}}\{x+iy:x,y\in\mathbb{R},\frac{x^{2}}{(1+\delta)^{2}}+\frac{y^{2}}{(1-\delta)^{2}}<1\},$ (2.1)
for some $\delta$ such that $10\leq\delta<$ 1. Note that any elliptical domain can be identified via an affine self-map of the plane with an elliptical domain of the form $G_{\delta}$ for some $\delta\in[0,1)$.
In this paper all Hilbert spaces are complex Hilbert spaces. For a complex Hilbert space $\mathcal{H}$ we denote by $\mathcal{B}(\mathcal{H})$ the space of bounded operators on $\mathcal{H}$. If $T\in\mathcal{B}(\mathcal{H})$, then $W(T)$, the numerical range of $T$, is defined by the formula
$W(T)=\{\langle Tx,x\rangle_{\mathcal{H}}\,:\,x\in\mathcal{H},\|x\|=1\}.$
The B. and F. Delyon family, $\mathcal{F}_{\mathrm{bfd}}(C)$, corresponding to an open bounded convex set $C$ in $\mathbb{C}$ is the class of operators $T$ such that the closure of the numerical range of $T$, $\overline{W(T)}$, is contained in $C$. By [18, Theorem 1.2-1], the spectrum $\sigma(T)$ of an operator $T$ is contained in $\overline{W(T)}$, and so, by the Riesz-Dunford functional calculus, $\varphi(T)$ is defined for all $\varphi\in{\rm Hol}(C)$ and $T\in\mathcal{F}_{\mathrm{bfd}}(C)$. Therefore, we may consider the calcular norm [^3] A calcular norm on a function space is a norm that is defined with the aid of the functional calculus. For more information on such norms the reader may consult [6, Chapter 9].
$\|\varphi\|_{\mathcal{F}_{\mathrm{bfd}}(C)}=\sup_{T\in\mathcal{F}_{\mathrm{bfd}}(C)}\|\varphi(T)\|,$ (2.2)
defined for $\varphi\in{\rm Hol}(C)$, and the associated Banach algebra
${\rm H}^{\infty}_{\mathrm{bfd}}(C)=\{\varphi\in{\rm Hol}(C)\,:\,\|\varphi\|_{\mathcal{F}_{\mathrm{bfd}}(C)}<\infty\}.$
In this paper the convex set $C$ will always be $G_{\delta}$, and so we abbreviate the notation to $\|\cdot\|_{\mathrm{bfd}}$ in place of $\|\cdot\|_{\mathcal{F}_{\mathrm{bfd}}(G_{\delta})}$. Thus
$\|\varphi\|_{\mathrm{bfd}}=\sup_{T\in\mathcal{F}_{\mathrm{bfd}}(G_{\delta})}\|\varphi(T)\|,$ (2.3)
defined for $\varphi\in{\rm Hol}(G_{\delta})$. In addition we introduce the bfd-Schur class, $\mathcal{S}_{\mathrm{bfd}}$, of functions on $G_{\delta}$, which is the set of functions $f\in{\rm Hol}(G_{\delta})$ such that $\|f\|_{\mathrm{bfd}}\leq 1.$ [^4] In the notations $\|\cdot\|_{\mathrm{bfd}}$ and $\mathcal{S}_{\mathrm{bfd}}$ we suppress dependence on the parameter $\delta$. The bfd-norm is named in recognition of a celebrated theorem [12] of the brothers B. and F. Delyon, which states that, if $p$ is a polynomial, $\mathcal{H}$ is a Hilbert space and $T\in\mathcal{B}(\mathcal{H})$ then
$\|p(T)\|\leq\kappa(W(T))\|p\|_{W(T)},$
where $\|\cdot\|_{W(T)}$ denotes the supremum norm on $W(T)$, and, for any bounded convex set $C$ in $\mathbb{C}$, $\kappa(C)$ is defined by
$\kappa(C)=3+\left(\frac{2\pi(\mathrm{diam}(C))^{2}}{\mathrm{area}(C)}\right)^{3}.$
Let us write
$K(\mathcal{F}_{\mathrm{bfd}}(C))=\sup_{\varphi\in{\rm Hol}(C):\|\varphi\|_{{\rm H}^{\infty}(C)}\leq 1}\|\varphi\|_{\mathcal{F}_{\mathrm{bfd}}(C)},$
and the Crouzeix universal constant
$K_{\mathrm{bfd}}=\sup\{K(\mathcal{F}_{\mathrm{bfd}}(C)):\ C\ \text{is a bounded convex set in }\ \mathbb{C}\}.$
In [9], Crouzeix proved $K_{\mathrm{bfd}}\leq 12$ and conjectured that $K_{\mathrm{bfd}}=2$. Subsequently Crouzeix and Palencia [10] proved that $K_{\mathrm{bfd}}\leq 1+\sqrt{2}$.
Let $\pi:R_{\delta}\to G_{\delta}$ be defined by $\pi(z)=z+\frac{\delta}{z}$, $z\in R_{\delta}$. Now observe that if $\varphi\in{\rm Hol}(G_{\delta})$ then we may define $\pi^{\sharp}(\varphi)\in{\rm Hol}(R_{\delta})$ by the formula
$\pi^{\sharp}(\varphi)(\lambda)=\varphi(\pi(\lambda))\qquad\text{ for all }\lambda\in R_{\delta}.$
We record the following simple fact from complex analysis without proof.
Lemma 2.4.
Let $\delta\in(0,1)$ and let $\psi\in{\rm Hol}(R_{\delta})$. Then $\psi\in\operatorname{ran}\pi^{\sharp}$ if and only if $\psi$ is symmetric with respect to the involution $\lambda\mapsto\delta/\lambda$ of $R_{\delta}$, that is, if and only if $\psi$ satisfies
$\psi(\delta/\lambda)=\psi(\lambda)$
for all $\lambda\in R_{\delta}$.
The following result, which is [4, Theorem 11.25], gives an intimate connection between the $\|\cdot\|_{\mathrm{dp}}$ and $\|\cdot\|_{\mathrm{bfd}}$ norms.
Theorem 2.5.
Let $\delta\in(0,1)$. The mapping $\pi^{\sharp}$ is an isometric isomorphism from ${\rm H}^{\infty}_{\mathrm{bfd}}(G_{\delta})$ onto the set of symmetric functions with respect to the involution $\lambda\mapsto\delta/\lambda$ in ${\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$, so that, for all $\varphi\in{\rm Hol}(G_{\delta})$,
$\|\varphi\|_{\mathrm{bfd}}=\|\varphi\circ\pi\|_{\mathrm{dp}}.$ (2.6)
Remark 2.7.
One can see that, for $\varphi\in{\rm Hol}(R_{\delta})$,
$\|\varphi\|_{\mathrm{dp}}$ $=\sup_{X\in\mathcal{F}_{\mathrm{dp}}(\delta)}\|\varphi(X)\|$
$\geq\sup_{X\in\mathcal{F}_{\mathrm{dp}}(\delta)\ \text{and}\ X\ \text{ is a scalar operator}\ }\|\varphi(X)\|$
$=\sup_{\lambda\in R_{\delta}}{|\varphi(\lambda)|}=\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}.$ (2.8)
Example 2.9.
Consider the function $f\in{\rm Hol}(R_{\delta})$ defined by $f(z)=z+\frac{\delta}{z}$. Then
$\|f\|_{\mathrm{dp}}=2\ \text{and}\ \|f\|_{{\rm H}^{\infty}(R_{\delta})}=1+\delta.$
Moreover, the Crouzeix universal constant $K_{\mathrm{bfd}}\geq 2$.
Proof.
If $\varphi(z)=z$ for $z\in G_{\delta}$ and $\pi:R_{\delta}\to G_{\delta}$ is defined by $\pi(z)=z+\frac{\delta}{z}$, then
$\varphi\circ\pi(z)=z+\frac{\delta}{z}=f(z)\ \text{ for}\ z\in R_{\delta}.$
By Theorem 2.5,
$\|\varphi\|_{\mathrm{bfd}}=\|\varphi\circ\pi\|_{\mathrm{dp}}.$
By [5, Example 4.26], $\|\varphi\|_{\mathrm{bfd}}=2$. Therefore,
$\|f\|_{\mathrm{dp}}=\|\varphi\circ\pi\|_{\mathrm{dp}}=\|\varphi\|_{\mathrm{bfd}}=2.$
Note that
$\|f\|_{{\rm H}^{\infty}(R_{\delta})}=\sup_{z\in R_{\delta}}\left|z+\frac{\delta}{z}\right|=1+\delta.$
Note that $\varphi(z)=z$ has bfd-norm equal to $2$ and sup norm on $G_{\delta}$ equal to $1+\delta$. Hence the Crouzeix universal constant $K_{\mathrm{bfd}}\geq 2$. ∎
Remark 2.10.
The inequality $K_{\mathrm{bfd}}\geq 2$ was proved by G. Tsikalas [22]. He considered the domain $A_{R}=\{1/R<|z|<R\}$ where $R>1,$ and the functions $g_{n}$ in ${\rm Hol}(A_{R})$ defined by
$g_{n}(z)=\frac{1}{R^{n}}\left(\frac{1}{z^{n}}+z^{n}\right),\ n=1,2,\dots.$
He showed that
$\limsup_{n\to\infty}\frac{\|g_{n}\|_{\mathrm{dp}}}{\|g_{n}\|_{{\rm H}^{\infty}(A_{R})}}\geq 2,$
from which it follows that $K_{\mathrm{bfd}}\geq 2$.
Proposition 2.11.
If $\varphi\in{\rm Hol}(\mathbb{D})$, then
$\|\varphi\|_{\mathrm{dp}}$ $=\sup_{X\in\mathcal{F}_{\mathrm{dp}}(\delta)}\|\varphi(X)\|$
$=\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}=\|\varphi\|_{{\rm H}^{\infty}(\mathbb{D})}.$ (2.12)
Proof.
By the definition of the dp-norm,
$\|\varphi\|_{\mathrm{dp}}$ $=\sup_{X\in\mathcal{F}_{\mathrm{dp}}(\delta)}\|\varphi(X)\|$
$\leq\sup_{\|X\|\leq 1}\|\varphi(X)\|$ by the definition of $\text{by the definition of }\mathcal{F}_{\mathrm{dp}}(\delta)$
$=\|\varphi\|_{{\rm H}^{\infty}(\mathbb{D})}.$ by von Neumann’s inequality (2.13)
By the Maximum principle, for $\varphi\in{\rm Hol}(\mathbb{D})$,
$\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}=\|\varphi\|_{{\rm H}^{\infty}(\mathbb{D})}.$ (2.14)
By inequality (2.7), $\|\varphi\|_{\mathrm{dp}}\geq\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}$ and, by inequality (2), $\|\varphi\|_{\mathrm{dp}}\leq\|\varphi\|_{{\rm H}^{\infty}(\mathbb{D})}$, and so
$\|\varphi\|_{\mathrm{dp}}=\|\varphi\|_{{\rm H}^{\infty}(\mathbb{D})}.$ (2.15)
Therefore, the equalities (2.11) hold. ∎
3. Models and realizations of holomorphic functions on $R_{\delta}$
In this section we review some known results on the function theory of holomorphic functions in the dp-norm on an annulas. The models and realizations of holomorphic functions $\varphi:R_{\delta}\to\mathbb{C}$ such that $\|\varphi\|_{\mathrm{dp}}\leq 1$ are presented in [6, Theorem 9.46]. The theorem states the following.
Theorem 3.1.
Let $\delta\in(0,1)$. Let $\varphi:R_{\delta}\to\mathbb{C}$ be holomorphic and satisfy $\|\varphi\|_{\mathrm{dp}}\leq 1$. There exists a $\mathrm{dp}$-model $(\mathcal{N},v)$ of $\varphi$ with parameter $\delta$, in the sense that there are Hilbert spaces $\mathcal{N}^{+},\mathcal{N}^{-}$ and an ordered pair $v=(v^{+},v^{-})$ of holomorphic functions, where $v^{+}:R_{\delta}\to\mathcal{N}^{+}$ and $v^{-}:R_{\delta}\to\mathcal{N}^{-}$ satisfy, for all $z,w\in R_{\delta}$,
$1-\overline{\varphi(w)}\varphi(z)=(1-\bar{w}z)\langle v^{+}(z),v^{+}(w)\rangle_{\mathcal{N}^{+}}+(\bar{w}z-\delta^{2})\langle v^{-}(z),v^{-}(w)\rangle_{\mathcal{N}^{-}}.$
Definition 3.2.
A positive semi-definite function on a set $X$ is a function $A:X\times X\to\mathbb{C}$ such that, for any positive integer $n$ and any points $x_{1},\dots,x_{n}\in X$, the $n\times n$ matrix $[A(x_{j},x_{i})]_{i,j=1}^{n}$ is positive semi-definite.
We shall write
$[A(x,y)]\geq 0,\ \text{for all}\ x,y\in X,$
to mean that $A$ is a positive semi-definite function on $X$.
Theorem 3.3.
$\varphi\in\mathcal{S}_{\mathrm{dp}}$ if and only if there exist a pair of positive semi-definite functions $A$ and $B$ on $R_{\delta}$ such that
$1-\overline{\varphi(\mu)}\varphi(\lambda)=(1-\bar{\mu}\lambda)A(\lambda,\mu)+(1-\overline{\frac{\delta}{\mu}}\frac{\delta}{\lambda})B(\lambda,\mu)$ (3.4)
for all $\lambda,\mu\in R_{\delta}$.
Proof.
For a proof see Definition 9.44 and Theorem 9.46 in [6]. ∎
Recall Moore’s Theorem [6, Theorem 2.5]: if $\Omega$ is a set and $A:\Omega\times\Omega\to\mathbb{C}$ is a function, then $A$ is a positive semi-definite function on $\Omega$ if and only if there exists a Hilbert space $\mathcal{M}$ and a function $u:\Omega\to\mathcal{M}$ satisfying
$A(\lambda,\mu)=\langle u(\lambda),u(\mu)\rangle_{\mathcal{M}}$ (3.5)
for all $\lambda,\mu\in\Omega$. Thus, if $A$ and $B$ are as in equation (3.4), we may choose Hilbert spaces $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ such that
$A(\lambda,\mu)=\langle u_{1}(\lambda),u_{1}(\mu)\rangle_{\mathcal{M}_{1}}\ \text{ and }\ B(\lambda,\mu)=\langle u_{2}(\lambda),u_{2}(\mu)\rangle_{\mathcal{M}_{2}}$
for all $\lambda,\mu\in R_{\delta}$. If we then let $\mathcal{M}=\mathcal{M}_{1}\oplus\mathcal{M}_{2}$ and define $E:R_{\delta}\to\mathcal{B}(\mathcal{M})$ and $u:R_{\delta}\to\mathcal{M}$ by the formulae
$E(\lambda)=\begin{bmatrix}\lambda&0\\ 0&\frac{\delta}{\lambda}\end{bmatrix}\ \text{ and }\ u(\lambda)=\begin{bmatrix}u_{1}(\lambda)\\ u_{2}(\lambda)\end{bmatrix},\qquad\text{for}\ \lambda\in R_{\delta},$ (3.6)
then the relation (3.4) becomes the formula
$1-\overline{\varphi(\mu)}\varphi(\lambda)=\Big{\langle}\,\big{(}1-E(\mu)^{*}E(\lambda)\big{)}\ u(\lambda)\,,u(\mu)\,\Big{\rangle}_{\mathcal{M}}\ \text{for}\ \lambda,\mu\in R_{\delta}.$ (3.7)
When $A$ is positive semi-definite, let us agree to say that $A$ has finite rank if $\mathcal{M}$ in the formula (3.5) can be chosen to have finite dimension. In this case, we may define $\operatorname{rank}(A)$ by setting
$\operatorname{rank}(A)=\dim\mathcal{M}$
where $\mathcal{M}$ satisfying (3.5) is chosen to have minimal dimension. [^5] Equivalently, $\{u(\lambda)\,:\,\lambda\in\Omega\}$ spans $\mathcal{M}$.
Definition 3.8.
If $\varphi\in{\rm Hol}(R_{\delta})$ we say $\varphi$ is of finite type if $\varphi\in\mathcal{S}_{\mathrm{dp}}$ and, in addition, $\varphi$ can be represented as in equation (3.4) where $A$ and $B$ have finite rank.
Theorem 3.9.
A realization formula. Let $\varphi\in\mathcal{S}_{\mathrm{dp}}(R_{\delta})$. If $(\mathcal{M},u)$ is a model for $\varphi$ then there exists a unitary operator $L\in\mathcal{B}(\mathbb{C}\oplus\mathcal{M})$ such that if we decompose $L$ as a block operator matrix
$L=\begin{bmatrix}a&1\otimes\beta\\ \gamma\otimes 1&D\end{bmatrix},$ (3.10)
where $a\in\mathbb{C}$, $\beta\in\mathcal{M}$, $\gamma\in\mathcal{M}$, and $D\in\mathcal{B}(\mathcal{M})$, then
$\varphi(\lambda)=a+\Big{\langle}\,E(\lambda)\big{(}1-DE(\lambda)\big{)}^{-1}\ \gamma\,,\beta\,\Big{\rangle}_{\mathcal{M}},\qquad\text{for all}\ \lambda\in R_{\delta}.$ (3.11)
Conversely, if $a\in\mathbb{C}$, $\beta\in\mathcal{M}$, $\gamma\in\mathcal{M}$, and $D\in\mathcal{B}(\mathcal{M})$ are such that $L$ as defined by equation (3.10) is unitary and if $\varphi$ is given by equation (3.11) and $u:R_{\delta}\to\mathcal{M}$ is defined by
$u(\lambda)=\big{(}1-DE(\lambda)\big{)}^{-1}\gamma,\qquad\text{for}\ \lambda\in R_{\delta},$ (3.12)
then $(\mathcal{M},u)$ is a model for $\varphi$.
Proof.
Let $(\mathcal{M},u)$ be a model for $\varphi$. As explained in Theorem 3.1, it means that there exist Hilbert spaces $\mathcal{N}^{+}$ and $\mathcal{N}^{-}$ and maps $v^{+}:R_{\delta}\to\mathcal{N}^{+},v^{-}:R_{\delta}\to\mathcal{N}^{-}$ such that, for all $\lambda,\mu\in R_{\delta}$,
$1-\overline{\varphi(\mu)}\varphi(\lambda)=(1-\bar{\mu}\lambda)\langle v^{+}(\lambda),v^{+}(\mu)\rangle_{\mathcal{N}^{+}}+(\bar{\mu}\lambda-\delta^{2})\langle v^{-}(\lambda),v^{-}(\mu)\rangle_{\mathcal{N}^{-}}.$
Reshuffle this relation to
$1+$ $\langle\lambda v^{+}(\lambda),\mu v^{+}(\mu)\rangle_{\mathcal{N}^{+}}+\langle\delta v^{-}(\lambda),\delta v^{-}(\mu)\rangle_{\mathcal{N}^{-}}$
$=\overline{\varphi(\mu)}\varphi(\lambda)+\langle v^{+}(\lambda),v^{+}(\mu)\rangle_{\mathcal{N}^{+}}+\langle\lambda v^{-}(\lambda),\mu v^{-}(\mu)\rangle_{\mathcal{N}^{-}},$
and notice that this equation amounts to saying that the following families of vectors in $\mathbb{C}\oplus\mathcal{N}$, where $\mathcal{N}\stackrel{{\scriptstyle\rm def}}{{=}}\mathcal{N}^{+}\oplus\mathcal{N}^{-}$,
$\begin{pmatrix}1\\ \lambda v^{+}(\lambda)\\ \delta v^{-}(\lambda)\end{pmatrix}_{\lambda\in R_{\delta}}\quad\text{and}\quad\begin{pmatrix}\varphi(\lambda)\\ v^{+}(\lambda)\\ \lambda v^{-}(\lambda)\end{pmatrix}_{\lambda\in R_{\delta}}$
have the same gramian. Let the closed linear spans of these two families be $\mathcal{X}$ and $\mathcal{Y}$ respectively. By the Lurking Isometry Lemma [6, Lemma 2.18] there exists a linear isometry $L:\mathcal{X}\to\mathcal{Y}$ such that
$L\begin{pmatrix}1\\ \lambda v^{+}(\lambda)\\ \delta v^{-}(\lambda)\end{pmatrix}=\begin{pmatrix}\varphi(\lambda)\\ v^{+}(\lambda)\\ \lambda v^{-}(\lambda)\end{pmatrix}$ (3.13)
for all $\lambda\in R_{\delta}$. Since both $\mathcal{X}$ and $\mathcal{Y}$ are subspaces of $\mathbb{C}\oplus\mathcal{N}$, we may extend $L$ (possibly after enlarging the space $\mathcal{N}$) to an isometry $L:\mathcal{N}\to\mathcal{N}$ (see the discussion in [6, Remark 2.31] for this step). Write $L$ as a block operator matrix
$L\sim\begin{pmatrix}a&1\otimes\beta\\ \gamma\otimes 1&D\end{pmatrix}$ (3.14)
with respect to the orthogonal decomposition $\mathbb{C}\oplus(\mathcal{N}^{+}\oplus\mathcal{N}^{-})$ of $\mathbb{C}\oplus\mathcal{N}$ and define a map $u:R_{\delta}\to\mathcal{N}$ by
$u(\lambda)=\begin{pmatrix}v^{+}(\lambda)\\ \lambda v^{-}(\lambda)\end{pmatrix}.$
Then equation (3.13) yields the relations
$a+\langle E(\lambda)u(\lambda),\beta\rangle_{\mathcal{N}}$ $=\varphi(\lambda)$ (3.15)
$\gamma+DE(\lambda)u(\lambda)$ $=u(\lambda),$ (3.16)
where $E(\lambda)$ is given by equation (3.6). Since $\|D\|\leq 1$ and $\|E(\lambda)\|=\max\{|\lambda|,\frac{\delta}{|\lambda|}\}<1$ for all $\lambda\in R_{\delta}$, it follows that $1-DE(\lambda)$ is invertible for $\lambda\in R_{\delta}$, and hence
$u(\lambda)$ $=(1-DE(\lambda))^{-1}\gamma,$
$\varphi(\lambda)$ $=a+\langle E(\lambda)(1-DE(\lambda))^{-1}\gamma,\beta\rangle$
for all $\lambda\in R_{\delta}$, which is the desired realization formula (3.11).
Conversely, suppose that $a,\beta,\gamma.D$ are such that $L$ given by equation (3.14) is an isometry on $\mathbb{C}\oplus\mathcal{N}$ and that $\varphi$ is the function on $R_{\delta}$ defined by equation (3.11). Since $1-DE(\lambda)$ is invertible for all $\lambda\in R_{\delta}$ we may define a mapping $u:R_{\delta}\to\mathcal{N}$ by equation (3.12). Then the equations (3.15) hold. They may be written in the form
$L\begin{bmatrix}1\\ E(\lambda)u(\lambda)\otimes 1\end{bmatrix}=\begin{bmatrix}\varphi(\lambda)\otimes 1\\ u(\lambda)\otimes 1\end{bmatrix}\ \text{for}\ \lambda\in R_{\delta}.$
Thus, for any $\mu\in R_{\delta}$,
$\begin{bmatrix}1&1\otimes E(\mu)u(\mu)\end{bmatrix}L^{*}=\begin{bmatrix}1\otimes\varphi(\mu)&1\otimes u(\mu)\end{bmatrix}.$
Multiply the last two displayed equations together and use the fact that $L^{*}L=1$ to infer that, for any $\lambda,\mu\in R_{\delta}$,
$\begin{bmatrix}1&1\otimes E(\mu)u(\mu)\end{bmatrix}\begin{bmatrix}1\\ E(\lambda)u(\lambda)\otimes 1\end{bmatrix}=\begin{bmatrix}1\otimes\varphi(\mu)&1\otimes u(\mu)\end{bmatrix}\begin{bmatrix}\varphi(\lambda)\otimes 1\\ u(\lambda)\otimes 1\end{bmatrix},$
which multiplies out to give the relation, for all $\lambda,\mu\in R_{\delta}$,
$1-\overline{\varphi(\mu)}\varphi(\lambda)=\langle(1-E(\mu)^{*}E(\lambda))u(\lambda),u(\mu)\rangle_{\mathcal{N}},$
that is, $(\mathcal{N},u)$ is a DP-model of $\varphi$. ∎
Let us recall the interpolation problem we posed in the Introduction.
Definition 3.17.
The DP Pick Problem. Given $n$ distinct points $\lambda_{1},\ldots,\lambda_{n}$ in $R_{\delta}$ and $z_{1},\ldots,z_{n}\in\mathbb{C}$, does there exist a function $\varphi\in{\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ with $\|\varphi\|_{\mathrm{dp}}\leq 1$ such that
$\varphi(\lambda_{j})=z_{j}\qquad\text{for}\ j=1,\ldots,n?$ (3.18)
We say the DP Pick Problem (3.17) is solvable if there exists $\varphi\in\mathcal{S}_{\mathrm{dp}}$ satisfying equations (3.18).
The following theorem, which is a Pick interpolation theorem in the dp norm, is [6, Theorem 9.55].
Theorem 3.19.
Let $\delta\in(0,1)$. Let $\lambda_{1},\dots,\lambda_{n}$ be distinct points in $R_{\delta}$ and let $z_{1},\dots,z_{n}$ be arbitrary complex numbers. There exists $f\in{\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ such that $\|f\|_{\mathrm{dp}}\leq 1$ and
$f(\lambda_{i})=z_{i}\;\;\text{for}\;\;i=1,\dots,n,$
if and only if there exist a pair of $n\times n$ positive semi-definite matrices $A=[a_{ij}]$ and $B=[b_{ij}]$ such that
$1-\overline{z_{i}}z_{j}=(1-\bar{\lambda}_{i}\lambda_{j})a_{ij}+(1-\frac{\delta^{2}}{\bar{\lambda}_{i}\lambda_{j}})b_{ij}$
for $i,j=1,\dots,n$.
We also assert a dual theorem in terms of “DP Szegő kernels”, which we discuss in the next section.
4. DP-Szegő kernels and normalized DP-Szegő kernels for the tuple $(\lambda_{1},\dots,\lambda_{n})$
In this section we follow Abrahamse’s idea of using families of kernels to solve Pick interpolation problems.
Definition 4.1.
We say that a kernel $k:R_{\delta}\times R_{\delta}$ is a DP Szegő kernel on $R_{\delta}$ if
$[(1-\bar{\mu}\lambda)k(\lambda,\mu)]\geq 0\ \ \text{ and }\ \ [(1-\frac{\delta}{\overline{\mu}}\frac{\delta}{\lambda})k(\lambda,\mu)]\geq 0,\qquad\text{for all}\;\lambda,\mu\in R_{\delta}.$ (4.2)
We let
[…]
Definition 4.3.
Let $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ be an $n$-tuple of distinct points in $R_{\delta}$. We denote by $\mathcal{F}_{\rm dp}(\delta,\lambda)$ the set of operators $T$ in the Douglas-Paulsen family $\mathcal{F}_{\mathrm{dp}}(\delta)$ corresponding to the annulus $R_{\delta}$ that act on an $n$-dimensional Hilbert space $\mathcal{H}_{T}$ and satisfy
$\sigma(T)=\{\lambda_{1},\ldots,\lambda_{n}\}.$
If $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$, then, as $\dim\mathcal{H}_{T}=n$ and $\sigma(T)$ consists of $n$ distinct points, $T$ is diagonalizable, that is, there exist $n$ linearly independent vectors $e_{1},\ldots,e_{n}\in\mathcal{H}_{T}$ such that
$Te_{j}=\lambda_{j}e_{j}\qquad\text{for}\ j=1,\ldots,n.$ (4.4)
Let $g$ denote the gramian of the vectors $e_{1},\ldots,e_{n}$, that is,
$g=[g_{ij}],\text{ where }g_{ij}=\langle e_{j},e_{i}\rangle\qquad\text{for}\ i,j=1,\ldots,n,$ (4.5)
Then, we shall prove in Proposition 4.9 that $g=[g_{ij}]$ is a positive definite $n\times n$ matrix such that
$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]\geq 0\ \ \text{ and }\ \ \left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]\geq 0.$ (4.6)
Definition 4.7.
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. We define $\mathcal{G}_{\mathrm{dp}}(\lambda)$ to be the set of positive definite $n\times n$ matrices $g=[g_{ij}]$ such that
$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]\geq 0\ \ \text{ and }\ \ \left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]\geq 0.$ (4.8)
We call $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ a DP-Szegő kernel for the $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$.
Proposition 4.9.
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. Let $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$. Then the gramian $g=[g_{ij}]$ of vectors $e_{1},\ldots,e_{n}$ that satisfy the equations (4.4) and (4.5) is a positive definite $n\times n$ matrix which belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$.
Proof.
By assumption $T$ is a Douglas-Paulsen operator with parameter $\delta$ that acts on an $n$-dimensional Hilbert space $\mathcal{H}_{T}$, $T$ has $n$ linearly independent eigenvectors $e_{1},\dots,e_{n}$ corresponding to the eigenvalues $\lambda_{1},\dots,\lambda_{n}$ respectively and $g_{ij}=\langle e_{j},e_{i}\rangle$ for $i,j=1,\dots,n$. By the definition of the Douglas-Paulsen class, $\|T\|\leq 1$ and $\|\delta/T\|\leq 1$, so that, for any vector $x=\sum_{j=1}^{n}x_{j}e_{j}$, we have
00 0 $\leq\|x\|^{2}-\|Tx\|^{2}$
$=\langle\sum_{j=1}^{n}x_{j}e_{j},\sum_{i=1}^{n}x_{i}e_{i}\rangle-\langle\sum_{j=1}^{n}x_{j}\lambda_{j}e_{j},\sum_{i=1}^{n}\lambda_{i}x_{i}e_{i}\rangle$
$=\sum_{i,j=1}^{n}\overline{x_{i}}\left((1-\overline{\lambda_{i}}\lambda_{j})g_{ij}\right)x_{j}$
$=\begin{bmatrix}\overline{x_{1}}&\dots&\overline{x_{n}}\end{bmatrix}\begin{bmatrix}(1-\overline{\lambda_{i}}\lambda_{j})g_{ij}\end{bmatrix}\begin{bmatrix}x_{1}\\ \dots\\ x_{n}\end{bmatrix}.$
Thus
$\big{[}(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}\big{]}_{i,j=1}^{n}\geq 0.$
Likewise, the relation $\|\frac{\delta}{T}x\|\leq\|x\|$ holds for any vector $x=\sum_{j=1}^{n}x_{j}e_{j}\in\mathcal{H}_{T}$. Therefore, we have
00 0 $\leq\|x\|^{2}-\left\|\frac{\delta}{T}x\right\|^{2}$
$=\langle\sum_{j=1}^{n}x_{j}e_{j},\sum_{i=1}^{n}x_{i}e_{i}\rangle-\langle\sum_{j=1}^{n}\frac{\delta}{\lambda_{j}}x_{j}e_{j},\sum_{i=1}^{n}\frac{\delta}{\lambda_{i}}x_{i}e_{i}\rangle$
$=\sum_{i,j=1}^{n}\overline{x_{i}}\left(\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right)x_{j}$
$=\begin{bmatrix}\overline{x_{1}}&\dots&\overline{x_{n}}\end{bmatrix}\begin{bmatrix}(1-{\frac{\delta}{\overline{\lambda_{i}}}}\frac{\delta}{\lambda_{j}})g_{ij}\end{bmatrix}\begin{bmatrix}x_{1}\\ \dots\\ x_{n}\end{bmatrix}.$
Thus
$\big{[}\left(1-{\frac{\delta}{\overline{\lambda_{i}}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\big{]}_{i,j=1}^{n}\geq 0.$
Therefore, $g=[g_{ij}]$ is a positive definite DP-Szegő kernel for the $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$. ∎
Let $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$, so that $g>0$. By Moore’s theorem [6, Theorem 2.5], $g$ is the gramian matrix of a basis $e_{1},\dots,e_{n}$ of an $n$-dimensional Hilbert space ${\mathcal{H}}$.
Proposition 4.10.
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. Let $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$. Let $g$ be the gramian matrix of a basis $e_{1},\dots,e_{n}$ of an $n$-dimensional Hilbert space ${\mathcal{H}}$. Define $T\in\mathcal{B}(\mathcal{H})$ by
$Te_{j}=\lambda_{j}e_{j},\ j=1,\dots,n.$ (4.11)
Then $T\in\mathcal{F}_{\mathrm{dp}}(\delta,\lambda)$.
Proof.
Let us show that $T$ is a Douglas-Paulsen operator. If $x=\sum_{j=1}^{n}\xi_{j}e_{j}\in\mathcal{H}$, $Tx=\sum_{j=1}^{n}\xi_{j}\lambda_{j}e_{j}$ and
$\|Tx\|^{2}$ $=\big{\langle}\,\sum_{j=1}^{n}\xi_{j}\lambda_{j}e_{j},\sum_{i=1}^{n}\xi_{i}\lambda_{i}e_{i}\,\big{\rangle}$
$=\sum_{j,i=1}^{n}\xi_{j}\lambda_{j}\bar{\xi}_{i}\bar{\lambda}_{i}\langle\,e_{j},e_{i}\,\rangle$
$=\sum_{j,i=1}^{n}\bar{\lambda}_{i}\lambda_{j}\xi_{j}\bar{\xi}_{i}\,g_{ij}.$ (4.12)
Hence
$\|x\|^{2}-\|Tx\|^{2}=\sum_{j,i=1}^{n}(1-\bar{\lambda}_{i}\lambda_{j})\,g_{ij}\xi_{j}\bar{\xi}_{i}.$
By hypothesis,
$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]\geq 0,$
and so $\|x\|^{2}-\|Tx\|^{2}\geq 0$. Thus $\|T\|\leq 1$. Similarly, using the hypothesis
$\left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]\geq 0,$
one can show that $\|\delta T^{-1}\|\leq 1$. Therefore $T$ is a Douglas-Paulsen operator. In addition, by the definition (4.11) of $T$ ,
$\sigma(T)=\{\lambda_{1},\ldots,\lambda_{n}\}\subset R_{\delta}.$
thus $T\in\mathcal{F}_{\mathrm{dp}}(\delta,\lambda)$. ∎
Proposition 4.14.
Let $\lambda_{1},\ldots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$ and $z_{1},\ldots,z_{n}\in\mathbb{C}$. If the DP Pick Problem 3.17 is solvable, then, for any positive definite $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$ (4.15)
Proof.
By assumption, the DP Pick Problem 3.17 is solvable, that is, there exist a function $\varphi\in{\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ with $\|\varphi\|_{\mathrm{dp}}\leq 1$ and satisfying
$\varphi(\lambda_{j})=z_{j},\qquad j=1,\ldots,n.$ (4.16)
Let $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$, and so $g>0$. By Moore’s theorem [6, Theorem 2.5], $g$ is the gramian matrix of a basis $e_{1},\dots,e_{n}$ of an $n$-dimensional Hilbert space ${\mathcal{H}}$. Define $T\in\mathcal{B}(\mathcal{H})$ by
$Te_{j}=\lambda_{j}e_{j},\ j=1,\dots,n.$ (4.17)
By Proposition 4.10, $T\in\mathcal{F}_{\mathrm{dp}}(\delta,\lambda)$. By assumption, $\varphi\in\mathcal{S}_{\mathrm{dp}}$, and so $\|\varphi(T)\|\leq 1$. For any $x=\sum_{j=1}^{n}\xi_{j}e_{j}\in\mathcal{H}$,
$\varphi(T)x$ $=\varphi(T)\sum_{j=1}^{n}\xi_{j}e_{j}$
$=\sum_{j=1}^{n}\xi_{j}\varphi(\lambda_{j})e_{j}=\sum_{j=1}^{n}\xi_{j}z_{j}e_{j}.$ (4.18)
Therefore, by equation (4), the condition $\|\varphi(T)\|\leq 1$ translates into
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$
Definition 4.19.
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. We say that a DP-Szegő kernel $[g_{ij}]\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ is normalized if $g_{ii}=1$ for $i=1,\dots,n$.
Let $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$ denote the set of normalized DP-Szegő kernels for the $n$-tuple $(\lambda_{1},\dots,\lambda_{n})$.
Remark 4.20.
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. Every DP-Szegő kernel $[g_{ij}]$ from $\mathcal{G}_{\mathrm{dp}}(\lambda)$ is diagonally congruent to a normalized DP-Szegő kernel.
Proof.
For any matrix $[g_{ij}]\in\mathcal{G}_{\mathrm{dp}}(\lambda)$, we can define a positive definite matrix $[h_{ij}]$ by
$h_{ii}=1\ \text{for}\ i=1,\dots,n\ \text{and}\ h_{ij}=c_{i}^{-1}g_{ij}c_{j}^{-1}\ \text{if}\ i\neq j$ (4.21)
where
$c_{i}=\sqrt{g_{ii}}\ \text{if}\ g_{ii}\neq 0\ \text{and}\ c_{i}=1\ \text{if}\ g_{ii}=0.$ (4.22)
Then $h_{ii}=1$ for each $i$, and
$\begin{bmatrix}h_{ij}\end{bmatrix}_{i,j=1}^{n}=C^{*}\begin{bmatrix}g_{ij}\end{bmatrix}_{i,j=1}^{n}C\ \text{where}\ C=\mathrm{diag}\{1/c_{1},\dots,1/c_{n}\}.$ (4.23)
On conjugating the inequalities (4.8) by the matrix $C$ we find that $[h_{ij}]$ belongs to $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. ∎
Proposition 4.24.
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. The set $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$ is compact in the topology of the space of $n\times n$ complex matrices. Moreover, for fixed target data $z_{1},\dots,z_{n}$,
$\begin{bmatrix}(1-\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}\geq 0\ \text{for all}\ g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ (4.25)
if and only if
$\begin{bmatrix}(1-\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}\geq 0\ \text{for all}\ g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda).$ (4.26)
Proof.
Consider any matrix $g=[g_{ij}]\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. Since $g$ is positive definite, the principal minor on rows $i$ and $j$ is non-negative, which is to say that $1-|g_{ij}|^{2}\geq 0$ for $i,j=1,\dots,n$. It follows that the operator norm $\|g\|\leq n$, and so $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$ is bounded. Let us prove that $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$ is sequentially compact.
Let $g^{\ell}$, $\ell=1,2,\dots$, be a sequence in $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. We claim that $(g^{\ell})_{\ell\geq 1}$ has a subsequence that converges to an element of $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. For each $\ell$, since $g^{\ell}$ is non-singular, by the definition of $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, we may pick a basis $e_{1}^{\ell},\dots,e_{n}^{\ell}$ of $\mathbb{C}^{n}$ such that $g^{\ell}$ is the gramian matrix of the basis $e_{1}^{\ell},\dots,e_{n}^{\ell}$, which is to say that
$g^{\ell}=[g_{ij}^{\ell}],\text{ where }g_{ij}^{\ell}=\langle e_{j}^{\ell},e_{i}^{\ell}\rangle\qquad\text{for}\ i,j=1,\ldots,n.$ (4.27)
Define $T^{\ell}\in\mathcal{B}(\mathbb{C}^{n})$ by
$T^{\ell}e_{j}^{\ell}=\lambda_{j}e_{j}^{\ell},\ j=1,\dots,n.$ (4.28)
Note that since $g^{\ell}$ is normalised, that is,
$g_{ii}^{\ell}=\langle e_{i}^{\ell},e_{i}^{\ell}\rangle=1\qquad\text{for}\ i=1,\ldots,n,$
and so $\|e_{i}^{\ell}\|=1$ for $i=1,\ldots,n$. Then, by Proposition 4.10,
$\sigma(T^{\ell})=\{\lambda_{1},\ldots,\lambda_{n}\}$
and $T^{\ell}\in\mathcal{F}_{\mathrm{dp}}(\delta,\lambda)$. By the compactness of the unit sphere in $\mathbb{C}^{n}$, we can choose a subsequence $(e_{i}^{\ell_{k}})_{k\geq 1}$ of $(e_{i}^{\ell})_{\ell\geq 1}$ such that $(e_{j}^{\ell_{k}})$ converges to a unit vector $v_{j}\in\mathbb{C}^{n}$ as $k\to\infty$ for $j=1,\dots,n$. By the compactness of the unit ball in $\mathcal{B}(\mathbb{C}^{n})$, by passing to a further subsequence $(e^{\ell_{k}})_{k\geq 1}$ of $(e^{\ell})_{\ell\geq 1}$ we can arrange also that $(T^{\ell_{k}})$ converges to a limit $T\in\mathcal{B}(\mathbb{C}^{n})$ as $k\to\infty$. In the relations
$T^{\ell_{k}}e_{j}^{\ell_{k}}=\lambda_{j}e_{j}^{\ell_{k}},\ j=1,\dots,n,$ (4.29)
let $k\to\infty$ to obtain
$Tv_{j}=\lambda_{j}v_{j}\qquad\text{and }\ \|v_{j}\|=1\ \qquad j=1,\dots,n.$ (4.30)
Thus
$\sigma(T^{\ell})=\{\lambda_{1},\ldots,\lambda_{n}\},$
the eigenvectors $v_{1},\dots,v_{n}$ of $T$ corresponding to the distinct eigenvalues $\lambda_{1},\ldots,\lambda_{n}$ are linearly independent and therefore span $\mathbb{C}^{n}$, and $T^{\ell}\in\mathcal{F}_{\mathrm{dp}}(\delta,\lambda)$. Let $g$ be the Gramian of the vectors $v_{1},\dots,v_{n}$ in $\mathbb{C}^{n}$: then $g$ is positive definite, and by Proposition 4.9, $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. We have
$g_{ij}=\langle v_{j},v_{i}\rangle=\lim_{k\to\infty}\langle v_{j}^{\ell_{k}},v_{i}^{\ell_{k}}\rangle=\lim_{k\to\infty}g_{ij}^{\ell_{k}}$
for $i,j=1,\dots,n$, and so $g^{\ell_{k}}\to g$ as $k\to\infty$. We have shown that $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$ is sequentially compact in the metrizable topology of $\mathcal{B}(\mathbb{C}^{n})$, hence it is compact.
To prove the “Moreover”, fix target data $z_{1},\dots,z_{n}$. Since $\mathcal{G}_{\mathrm{dp}}(\lambda)\supset\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, trivially statement (4.25) implies statement (4.26). Conversely, suppose statement (4.26) holds and consider any kernel $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$. Define matrices $h=[h_{ij}]$ and $C$ by the relations (4.21), (4.22) and (4.23). Then $h\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, and so, by assumption,
$[(1-\bar{z}_{i}z_{j})h_{ij}]\geq 0$ (4.31)
Conjugate this matrix inequality by $\mathrm{diag}\{c_{1},\dots,c_{n}\}$ to obtain the relation (4.25). Thus the relation (4.26) implies the relation (4.25). ∎
Definition 4.32.
Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. Let $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ be a DP Szegő kernel for the $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$. The matrices
$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]_{i,j=1}^{n}\ \ \text{ and }\ \ \left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]_{i,j=1}^{n}$ (4.33)
are said to be the defect matrices of $g$.
Say that a DP-Szegő kernel $g$ on $R_{\delta}$ is reducible if there exist DP-Szegő kernels $h$ and $k$ on $R_{\delta}$ such that $g=h+k$ and neither $h$ nor $k$ is diagonally congruent to $g$. Here two kernels $g$ and $h$ on $R_{\delta}$ are said to be diagonally congruent if there exists a function $c:R_{\delta}\to\mathbb{C}\setminus\{0\}$ such that, for all $\lambda,\mu\in R_{\delta}$, $h(\lambda,\mu)=c(\lambda)g(\lambda,\mu)\overline{c(\mu)}$. A DP-Szegő kernel is irreducible if it is not reducible. Clearly, if DP Pick data $\lambda_{j}\mapsto z_{j},j=1,\dots,n$, are such that
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0\ \text{and}\ [1-\frac{\delta^{2}}{\overline{z_{i}}z_{j}}g_{ij}]\geq 0$
for all irreducible DP Szegő kernels $g$ then the same inequality holds for all DP Szegő kernels, and consequently the DP pick interpolation problem is solvable. Since the class of irreducible DP Szegő kernels is likely to be much smaller than the class of all DP Szegő kernels, it would be valuable to identify the irreducible DP Szegő kernels on $R_{\delta}$.
5. The DP Pick problem and DP-Szegő kernels
Definition 5.1.
Let $H_{n}$ be the real linear space of Hermitian matrices in $\mathbb{C}^{n\times n}$. A subset $P$ of $H_{n}$ is called a cone if the following conditions are satisfied: (i) $P+P\subseteq P$, (ii) $P\cap(-P)=\{0\}$ and (iii) $\alpha P\subseteq P$ whenever $\alpha\in\mathbb{R}$ and $\alpha\geq 0$.
Theorem 5.2.
Let $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$ be distinct and let $z_{1},\dots,z_{n}\in\mathbb{C}$. There exists $\varphi\in\mathcal{S}_{\mathrm{dp}}$ such that
$\varphi(\lambda_{j})=z_{j}\qquad\text{for}\ j=1,\ldots,n,$
if and only if, for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$ (5.3)
Proof.
Implication $\Rightarrow$ follows from Proposition 4.14.
To prove $\Leftarrow$, suppose that
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0$ (5.4)
for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$.
By Theorem 3.19, to show that the DP Pick Problem (3.17) is solvable it suffices to prove that there exist a pair of $n\times n$ positive semi-definite matrices $A=[a_{ij}]$ and $B=[b_{ij}]$ such that
$1-\overline{z_{i}}z_{j}=(1-\bar{\lambda}_{i}\lambda_{j})a_{ij}+(\bar{\lambda}_{i}\lambda_{j}-\delta^{2})b_{ij}$
for all $i,j=1,\dots,n$. Let $H_{n}$ be the real linear space of Hermitian matrices in $\mathbb{C}^{n\times n}$, and let
${\mathcal{C}}=\left\{[(1-\bar{\lambda}_{i}\lambda_{j})a_{ij}]_{i,j=1}^{n}++\left[\left(1-\frac{\delta^{2}}{\bar{\lambda}_{i}\lambda_{j}}\right)b_{ij}\right]_{i,j=1}^{n}:\,[a_{ij}]_{i,j=1}^{n}\geq 0\ \text{and}\ [b_{ij}]_{i,j=1}^{n}\geq 0\right\}\subset H_{n}.$ (5.5)
The subset ${\mathcal{C}}$ is a closed convex cone in $H_{n}$.
Note that every $n\times n$ positive semi-definite matrix $[a_{ij}]_{i,j=1}^{n}$ belongs to ${\mathcal{C}}$. By the positivity of Szegő kernel $\left[\frac{1}{1-\bar{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{n}$, the $n\times n$ matrix of the form
$\left[\frac{a_{ij}}{1-\bar{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{n}$
is also positive semi-definite. In the definition of ${\mathcal{C}}$ (5.5) we may replace $[a_{ij}]_{i,j=1}^{n}$ by $\left[\frac{a_{ij}}{1-\bar{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{n}$ and $[b_{ij}]_{i,j=1}^{n}$ by the zero matrix, to deduce that $[a_{ij}]_{i,j=1}^{n}$ belongs to ${\mathcal{C}}$.
By the Hahn-Banach theorem, to show that $[1-\bar{z}_{i}z_{j}]_{i,j=1}^{n}$ belongs to ${\mathcal{C}}$ it suffices to prove that, for every real linear functional ${\mathcal{L}}$ on $H_{n}$, ${\mathcal{L}}\geq 0$ on ${\mathcal{C}}$ implies ${\mathcal{L}}([1-\bar{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0$.
Extend ${\mathcal{L}}$ to a complex linear functional $\tilde{\mathcal{L}}$ on $\mathbb{C}^{n\times n}$ by
~$\tilde{\mathcal{L}}(X+{\mathrm{i}}\,Y)={\mathcal{L}}(X)+{\mathrm{i}}\,{\mathcal{L}}(Y)$
for $X,Y\in H_{n}$. Now define a pre-inner product $\langle\cdot,\cdot\rangle_{L}$ on $\mathbb{C}^{n}$ by
$\langle\,c,d\,\rangle_{L}=\tilde{\mathcal{L}}(c\otimes d)$
for $c,d\in\mathbb{C}^{n}$. Here $c\otimes d\in\mathbb{C}^{n\times n}$, defined by
$(c\otimes d)(x)=\langle\,x,d\,\rangle_{\mathbb{C}^{n}}c\;\;\;\text{for all}\ x\in\mathbb{C}^{n}.$
Note that, for any $c\in\mathbb{C}^{n}$,
$\langle\,c,c\,\rangle_{L}=\tilde{\mathcal{L}}(c\otimes c)={\mathcal{L}}(c\otimes c)\geq 0.$
Let
${\mathcal{N}}=\{x\in\mathbb{C}^{n}:\,\langle\,x,x\,\rangle_{L}=0\}.$
Then ${\mathcal{N}}$ is a subspace of $\mathbb{C}^{n}$, and $\langle\cdot,\cdot\rangle_{L}$ induces an inner product on $\mathbb{C}^{n}/{\mathcal{N}}$.
Let $e_{1},\dots,e_{n}$ be the standard basis of $\mathbb{C}^{n}$ and let $T\in\mathcal{B}(\mathbb{C}^{n})$ defined by
$Te_{j}=\lambda_{j}e_{j},\ j=1,\dots,n.$ (5.6)
For $x=\sum_{j=1}^{n}\xi_{j}e_{j}\in\mathbb{C}^{n}$, we have
$\langle\,x,x\,\rangle_{L}-\langle\,Tx,Tx\,\rangle_{L}$ =~$=\tilde{\mathcal{L}}(x\otimes x)-\tilde{\mathcal{L}}(Tx\otimes Tx)$
=~$=\tilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}e_{i}\right)-\tilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}\lambda_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}\lambda_{i}e_{i}\right)$
=~$=\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\bar{\xi}_{i}e_{j}\,\otimes e_{i}\right)-\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\lambda_{j}\bar{\xi}_{i}\bar{\lambda}_{i}e_{j}\,\otimes e_{i}\right)$
=~$=\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}(1-\bar{\lambda}_{i}\lambda_{j})\bar{\xi}_{i}\xi_{j}e_{j}\,\otimes e_{i}\right)$
=~$=\tilde{\mathcal{L}}\left[(1-\bar{\lambda}_{i}\lambda_{j})\bar{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}$
$={\mathcal{L}}\left[(1-\bar{\lambda}_{i}\lambda_{j})\bar{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}\geq 0\;\;\;\text{since}\;\;\mathcal{L}\geq 0\ \text{on}\ {\mathcal{C}}.$ (5.7)
Thus
$\langle\,x,x\,\rangle_{L}-\langle\,Tx,Tx\,\rangle_{L}\geq 0,$
and so $x\in{\mathcal{N}}$ implies that $Tx\in{\mathcal{N}}$. Hence $T$ induces an operator $\tilde{T}$ on $\mathbb{C}^{n}/{\mathcal{N}}$ by
$\tilde{T}(x+{\mathcal{N}})=Tx+{\mathcal{N}},$
and $\|\tilde{T}(x+{\mathcal{N}})\|^{2}\leq\|x+{\mathcal{N}}\|^{2}$ for all $(x+{\mathcal{N}})\in\mathbb{C}^{n}/{\mathcal{N}}$.
Similarly, let us prove that $\|\delta\tilde{T}^{-1}\|\leq 1$. For $x=\sum_{j=1}^{n}\xi_{j}e_{j}\in\mathbb{C}^{n}$, we have
$\langle\,x,x\,\rangle_{L}-\langle\,\delta T^{-1}x,\delta T^{-1}x\,\rangle_{L}$ =~$=\tilde{\mathcal{L}}(x\otimes x)-\tilde{\mathcal{L}}(\delta T^{-1}x\otimes\delta T^{-1}x)$
=~$=\tilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}e_{i}\right)-\tilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}\frac{\delta}{\lambda_{j}}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}\frac{\delta}{\lambda_{i}}e_{i}\right)$
=~$=\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\bar{\xi}_{i}e_{j}\,\otimes e_{i}\right)-\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\frac{\delta}{\lambda_{j}}\bar{\xi}_{i}\frac{\delta}{\overline{\lambda_{i}}}e_{j}\,\otimes e_{i}\right)$
=~$=\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)\bar{\xi}_{i}\xi_{j}e_{j}\,\otimes e_{i}\right)$
=~$=\tilde{\mathcal{L}}\left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)\bar{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}$
$={\mathcal{L}}\left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)\bar{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}\geq 0\;\;\;\text{since}\;\;\mathcal{L}\geq 0\ \text{on}\ {\mathcal{C}}.$ (5.8)
Thus
$\langle\,x,x\,\rangle_{L}-\langle\,\delta T^{-1}x,\delta T^{-1}x\,\rangle_{L}\geq 0,$
and so $x\in{\mathcal{N}}$ implies that $\delta T^{-1}x\in{\mathcal{N}}$.
Therefore, $T$ is a Douglas-Paulsen operator. Since the eigenvalues of $T$, $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$, $\sigma(T)\subseteq R_{\delta}$, the operator $T\in\mathcal{F}_{\mathrm{dp}}(\delta)$. Therefore, by Proposition 4.9, $[\langle\,e_{j},e_{i}\,\rangle_{L}]_{i,j=1}^{n}$ is a DP Szegő kernel for the $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$.
By supposition,
$[(1-\bar{z}_{i}z_{j})\langle\,e_{j},e_{i}\,\rangle_{L}]_{i,j=1}^{n}\geq 0.$
Choose a polynomial $p$ such that $p(\lambda_{i})=z_{i}$, $i=1,\dots,n$. Then ${}p(\tilde{T})e_{i}=z_{i}e_{i}$, $i=1,\dots,n$, and so $\|p(\tilde{T})\|\leq 1$. Choose $c=\begin{bmatrix}1\\ \dots\\ 1\end{bmatrix}\in\mathbb{C}^{n}$. Then
$\langle\,(1-p(\tilde{T})^{*}p(\tilde{T}))c,c\,\rangle_{L}\geq 0,$
that is,
${\mathcal{L}}([1-\bar{z}_{i}z_{j}]_{i,j=1}^{n}*cc^{*})\geq 0,\ \text{and so }\ {\mathcal{L}}([1-\bar{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0,$
where * denotes the Schur product of matrices.
Thus, for every real linear functional ${\mathcal{L}}$ on $H_{n}$ such that ${\mathcal{L}}\geq 0$ on ${\mathcal{C}}$ we have
${\mathcal{L}}([1-\bar{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0.$
Hence $[1-\bar{z}_{i}z_{j}]_{i,j=1}^{n}$ belongs to ${\mathcal{C}}$. ∎
Theorem 5.9.
Let $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$ be distinct and let $z_{1},\dots,z_{n}\in\mathbb{C}$. If the DP Pick problem
$\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,\ldots,n$
is solvable, then there exists a rational function $\varphi\in\mathcal{S}_{\mathrm{dp}}$ which satisfies the equations
$\varphi(\lambda_{j})=z_{j}\qquad\text{for}\ j=1,\ldots,n,$ (5.10)
and has a model $(\mathcal{M},u)$, with $u:R_{\delta}\to\mathcal{M}$ holomorphic, so that
$1-\overline{\varphi(\mu)}\varphi(\lambda)=\Big{\langle}\,\big{(}1-E(\mu)^{*}E(\lambda)\big{)}\ u(\lambda)\,,u(\mu)\,\Big{\rangle}_{\mathcal{M}}\ \text{for}\ \lambda,\mu\in R_{\delta},$ (5.11)
where $\mathcal{M}$ can be written as $\mathcal{M}=\mathcal{M}_{1}\oplus\mathcal{M}_{2}$, $\dim\mathcal{M}\leq 2n$ and $E:R_{\delta}\to\mathcal{B}(\mathcal{M})$ is defined by the formula
$E(\lambda)=\begin{bmatrix}\lambda&0\\ 0&\frac{\delta}{\lambda}\end{bmatrix},\ \qquad\text{for}\ \lambda\in R_{\delta},$ (5.12)
with respect to this orthogonal decomposition of $\mathcal{M}$.
Proof.
Suppose that
$\lambda_{j}\mapsto z_{j}\ \text{for}\ j=1,\dots,n$
is a solvable DP-Pick problem. By Theorem 3.19, there exist positive semi-definite $n\times n$ matrices $a=\begin{bmatrix}a_{ij}\end{bmatrix}$ and $b=\begin{bmatrix}b_{ij}\end{bmatrix}$ such that
$1-\bar{z}_{i}z_{j}=(1-\bar{\lambda}_{i}\lambda_{j})a_{ij}+(1-\frac{\delta^{2}}{\overline{\lambda_{i}}\lambda_{j}})b_{ij}\quad\text{for}\ i,j=1,\dots,n.$ (5.13)
Let the ranks of the matrices $a,b$ be $r_{1},r_{2}$ respectively, so that $r_{1}\leq n,r_{2}\leq n$. Then there exist vectors $x_{1},\dots,x_{n}\in\mathbb{C}^{r_{1}},y_{1},\dots,y_{n}\in\mathbb{C}^{r_{2}}$ such that
$a_{ij}=\langle x_{j},x_{i}\rangle_{\mathbb{C}^{r_{1}}}\ \text{and}\ b_{ij}=\langle y_{j},y_{i}\rangle_{\mathbb{C}^{r_{2}}}\quad\text{for}\ i,j=1,\dots,n.$
Substituting these relations into the equations (5.13) and re-arranging, we obtain the relations
$1+\langle\lambda_{j}x_{j},\lambda_{i}x_{i}\rangle_{\mathbb{C}^{r_{1}}}+\langle\frac{\delta}{\lambda_{j}}y_{j},\frac{\delta}{\lambda_{i}}y_{i}\rangle_{\mathbb{C}^{r_{2}}}=\bar{z}_{i}z_{j}+\langle x_{j},x_{i}\rangle_{\mathbb{C}^{r_{1}}}+\langle y_{j},y_{i}\rangle_{\mathbb{C}^{r_{2}}}\ \text{for}\ i,j=1,\dots,n.$
These equations can in turn be expressed by saying that the families of vectors
$\begin{pmatrix}1\\ \lambda_{j}x_{j}\\ \frac{\delta}{\lambda_{j}}y_{j}\end{pmatrix}_{j=1,\dots,n}\ \text{and}\ \begin{pmatrix}z_{j}\\ x_{j}\\ y_{j}\end{pmatrix}_{j=1,\dots,n}$
in $\mathbb{C}\oplus\mathbb{C}^{r_{1}}\oplus\mathbb{C}^{r_{2}}$ have the same gramians. It follows from the “lurking isometry lemma” [6, Lemma 2.18] that there exists an isometry $L\in\mathcal{B}(\mathbb{C}\oplus\mathbb{C}^{r_{1}}\oplus\mathbb{C}^{r_{2}})$ such that
$L\begin{pmatrix}1\\ \lambda_{j}x_{j}\\ \frac{\delta}{\lambda_{j}}y_{j}\end{pmatrix}=\begin{pmatrix}z_{j}\\ x_{j}\\ y_{j}\end{pmatrix}\ \text{for}\ j=1,\dots,n.$ (5.14)
Express $L$ by an operator matrix with respect to the orthogonal decomposition $\mathbb{C}\oplus\left(\mathbb{C}^{r_{1}}\oplus\mathbb{C}^{r_{2}}\right)$:
$L\sim\begin{bmatrix}a&1\otimes\beta\\ \gamma\otimes 1&D\end{bmatrix},$
where $a\in\mathbb{C},\ \beta,\gamma\in\mathbb{C}^{r_{1}}\oplus\mathbb{C}^{r_{2}}$ and $D\in\mathcal{B}(\mathbb{C}^{r_{1}}\oplus\mathbb{C}^{r_{2}})$. In terms of these variables and our previous notation
$E(\lambda)\stackrel{{\scriptstyle\rm def}}{{=}}\begin{bmatrix}\lambda&0\\ 0&\frac{\delta}{\lambda}\end{bmatrix}:\mathbb{C}^{r_{1}}\oplus\mathbb{C}^{r_{2}}\to\mathbb{C}^{r_{1}}\oplus\mathbb{C}^{r_{2}}\ \text{for}\ \lambda\in R_{\delta},$
equation (5.14) can be written
$a+\langle E(\lambda_{j})\begin{pmatrix}x_{j}\\ y_{j}\end{pmatrix},\beta\rangle_{\mathbb{C}^{r_{1}}\oplus\mathbb{C}^{r_{2}}}$ $=z_{j}$
$\gamma+DE(\lambda_{j})\begin{pmatrix}x_{j}\\ y_{j}\end{pmatrix}$ $=\begin{pmatrix}x_{j}\\ y_{j}\end{pmatrix}$ (5.15)
for $j=1,\dots,n$. Observe that, for any $\lambda\in R_{\delta}$, $\|E(\lambda)\|<1$. As also $\|D\|\leq 1$ (since $L$ is an isometry), $1-DE(\lambda_{j})$ is invertible for each $j$. The equations (5) can therefore be solved to give
$\begin{pmatrix}x_{j}\\ y_{j}\end{pmatrix}$ $=(1-DE(\lambda_{j}))^{-1}\gamma$
$z_{j}$ $=a+\langle E(\lambda_{j})(1-DE(\lambda_{j}))^{-1}\gamma,\beta\rangle.$ (5.16)
Now define $\varphi\in{\rm Hol}(R_{\delta})$ by
$\varphi(\lambda)=a+\langle E(\lambda)(1-DE(\lambda))^{-1}\gamma,\beta\rangle_{\mathbb{C}^{r_{1}}\oplus\mathbb{C}^{r_{2}}},\ \text{for}\ \lambda\in R_{\delta}.$ (5.17)
By equation (5), $\varphi(\lambda_{j})=z_{j}$ for $j=1,\dots,n$, and by [6, Theorem 9.54], $\varphi\in\mathcal{S}_{\mathrm{dp}}$, while equation (5.17) constitutes a DP-realization for $\varphi$. By Cramer’s rule for an invertible matrix, the function $\varphi$ defined by equation (5.17) is a rational function. Accordingly, by Theorem 3.9, if we set $\mathcal{M}=\mathbb{C}^{r_{1}}\oplus\mathbb{C}^{r_{2}}$ and define a holomorphic function $u:R_{\delta}\to\mathcal{M}$ by $u(\lambda)=(1-E(\lambda)D)^{-1}\gamma$, for $\lambda\in R_{\delta}$, then $(\mathcal{M},u)$ as in equation (5.11) is a DP-model for $\varphi$, while clearly $\dim\mathcal{M}=r_{1}+r_{2}\leq 2n$. ∎
Remark 5.18.
Solvable Pick data on $\mathbb{D}$ are also solvable as DP Pick data. Let $\lambda_{1},\ldots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$ and $z_{1},\ldots,z_{n}\in\mathbb{C}$. Suppose the Pick interpolation problem on the open unit disc $\mathbb{D}$
$\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,\ldots,n$
is solvable. Then the DP Pick Problem
$\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,\ldots,n$
is also solvable.
Proof.
By the assumption, there exists a holomorphic function $\varphi:\mathbb{D}\to\mathbb{C}$ such that $\varphi(\lambda_{j})=z_{j}$ for $j=1,\dots,n$ and $\|\varphi\|_{{\rm H}^{\infty}(\mathbb{D})}\leq 1$. By Proposition 2.11,
$\|\varphi|R_{\delta}\|_{\mathrm{dp}}=\|\varphi\|_{{\rm H}^{\infty}(\mathbb{D})}\leq 1,$ (5.19)
and so the restriction of $\varphi$ to $R_{\delta}$ is in $\mathcal{S}_{\mathrm{dp}}$, which is to say that the corresponding DP Pick problem is solvable. ∎
Example 5.20.
A solvable DP Pick data-set which is not a solvable Pick data-set on $\mathbb{D}$.
Let $\delta\in(0,\tfrac{1}{2})$ and consider the $2$ distinct points $\lambda_{1}=\tfrac{1}{2},\lambda_{2}=-\tfrac{1}{2}$ in $R_{\delta}$. Recall that in Example 2.9 we showed that the function $\varphi\in{\rm Hol}(R_{\delta})$, $\varphi(\lambda)=\tfrac{1}{2}(\lambda+\frac{\delta}{\lambda})$ satisfies $\|\varphi\|_{\mathrm{dp}}=1$. Let $z_{1}=\varphi(\lambda_{1})=\delta+\frac{1}{4}$, $z_{2}=\varphi(\lambda_{2})=-(\delta+\frac{1}{4})$. Thus the DP Pick Problem
$\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,2,$
is solvable by the function $\varphi(\lambda)=\tfrac{1}{2}(\lambda+\frac{\delta}{\lambda})$.
As to the Pick interpolation problem on the open unit disc $\mathbb{D}$
$\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,2,$
we have $3$ cases:
  1. (i)
    for $\delta\in(0,\frac{1}{4})$, the Pick interpolation problem
    $\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,2,$
    is solvable;
  2. (ii)
    for $\delta=\frac{1}{4}$, the Pick interpolation problem
    $\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,2,$
    on the open unit disc $\mathbb{D}$ is extremally solvable and has the unique solution $f(\lambda)=\lambda$;
  3. (iii)
    for $\delta\in(\frac{1}{4},\frac{1}{2})$, the Pick interpolation problem
    $\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,2,$
    is not solvable on $\mathbb{D}$.
Proof.
To prove (i)-(iii) on the solvability of the Pick interpolation problem on the open unit disc $\mathbb{D}$
$\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,2,$
we consider the appropriate Pick matrix, which here is
$P(\delta)=\displaystyle\begin{bmatrix}\frac{1-\overline{z_{i}}z_{j}}{1-\overline{\lambda_{i}}\lambda_{j}}\end{bmatrix}_{i,j=1}^{2}.$
That is,
$P(\delta)=\begin{bmatrix}\frac{1-(\delta+\frac{1}{4})^{2}}{1-\frac{1}{4}}&\frac{1+(\delta+\frac{1}{4})^{2}}{1+\frac{1}{4}}\\ \\ \frac{1+(\delta+\frac{1}{4})^{2}}{1+\frac{1}{4}}&\frac{1-(\delta+\frac{1}{4})^{2}}{1-\frac{1}{4}}.\\ \end{bmatrix}$ (5.21)
It is clear that, for $\delta\in(0,\tfrac{1}{2})$,
$P(\delta)_{11}=\frac{1-(\delta+\frac{1}{4})^{2}}{1-\frac{1}{4}}>0.$
A little calculation shows that the determinant of the Pick matrix
$\det P(\delta)=\frac{16^{2}}{15^{2}}\left\{\delta^{2}+\tfrac{1}{2}\delta-\frac{3}{16}\right\}\left\{\delta^{2}+\tfrac{1}{2}\delta-\frac{63}{16}\right\},$
from which one can deduce that
  1. (i)
    when $\delta\in(0,\frac{1}{4})$, $\det P(\delta)>0$ and so $P(\delta)>0$;
  2. (ii)
    $\det P(\delta)=0$, when $\delta=\frac{1}{4}$; and
  3. (iii)
    $\det P(\delta)<0$, when $\delta\in(\frac{1}{4},\frac{1}{2})$.
Therefore the Pick matrix $P(\delta)$ is not positive when $\delta\in(\frac{1}{4},\frac{1}{2})$. Thus, by Pick’s theorem, for $\delta\in(\frac{1}{4},\frac{1}{2})$, the Pick interpolation problem
$\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,2,$
is not solvable, while, for $\delta=\tfrac{1}{4}$, the Pick interpolation problem is uniquely solvable, and one sees by inspection that the unique solution is the function $f(\lambda)=\lambda$.
Example 5.22.
A solvable DP-Pick problem on $R_{\delta}$ which is not solvable as a Pick problem on $\mathbb{D}$. Let $\delta\in(0,1)$ and let $\lambda_{1}=\frac{\delta+\sqrt{\delta}}{2}$, $\lambda_{2}=-\lambda_{1}$. We have $10<\delta<\frac{\delta+\sqrt{\delta}}{2}<\sqrt{\delta}<$ 1, so that $\lambda_{1},\lambda_{2}\in R_{\delta}$. Recall from Example 2.9 that the function $\varphi(z)=\tfrac{1}{2}(z+\frac{\delta}{z})$ on $R_{\delta}$ satisfies $\|\varphi\|_{\mathrm{dp}}=1$. Consider the DP-Pick problem
$\lambda_{i}\mapsto z_{i}\stackrel{{\scriptstyle\rm def}}{{=}}\varphi(\lambda_{i}),i=1,2.$ (5.23)
Clearly this is a solvable DP-Pick problem, with solution $\varphi$. However, the Pick problem with the same data $\lambda_{j}\mapsto z_{j},j=1,2$, is not solvable. Indeed, the Pick matrix for the problem (5.23) is
$P=\begin{bmatrix}\frac{1-|z_{1}|^{2}}{1-|\lambda_{1}|^{2}}&\frac{1+|z_{1}|^{2}}{1+|\lambda_{1}|^{2}}\\ \frac{1+|z_{1}|^{2}}{1+|\lambda_{1}|^{2}}&\frac{1-|z_{1}|^{2}}{1-|\lambda_{1}|^{2}}\end{bmatrix}.$
Thus
$\det P$ $=\left(\frac{1-|z_{1}|^{2}}{1-|\lambda_{1}|^{2}}\right)^{2}-\left(\frac{1+|z_{1}|^{2}}{1+|\lambda_{1}|^{2}}\right)^{2}$
$=D_{1}D_{2}$
where
$D_{1}$ $=\frac{1-|z_{1}|^{2}}{1-|\lambda_{1}|^{2}}-\frac{1+|z_{1}|^{2}}{1+|\lambda_{1}|^{2}}$
$=\frac{2(|\lambda_{1}|^{2}-|z_{1}|^{2})}{1-|\lambda_{1}|^{4}},$
$D_{2}$ $=\frac{1-|z_{1}|^{2}}{1-|\lambda_{1}|^{2}}+\frac{1+|z_{1}|^{2}}{1+|\lambda_{1}|^{2}}$
$=\frac{2(1-|z_{1}\lambda_{1}|^{2})}{1-|\lambda_{1}|^{4}}.$
Now
$z_{1}$ $=\frac{1}{2}\left(\lambda_{1}+\frac{\delta}{\lambda_{1}}\right)=\frac{1}{2}\left(\frac{\delta+\sqrt{\delta}}{2}+\frac{2\delta}{\delta+\sqrt{\delta}}\right)$
$=\frac{\sqrt{\delta}}{2}\left(\frac{1+\sqrt{\delta}}{2}+\frac{2}{1+\sqrt{\delta}}\right)=\frac{\sqrt{\delta}(5+2\sqrt{\delta}+\delta)}{4(1+\sqrt{\delta})},$
and
$0<z_{1}\lambda_{1}$ $=\frac{\sqrt{\delta}(5+2\sqrt{\delta}+\delta)}{4(1+\sqrt{\delta})}.\frac{\sqrt{\delta}(1+\sqrt{\delta})}{2}=\frac{\delta(5+2\sqrt{\delta}+\delta)}{8}$
$<1.$
Thus $D_{2}>0$, and moreover
$|\lambda_{1}|-|z_{1}|$ $=\lambda_{1}-\tfrac{1}{2}(\lambda_{1}+\frac{\delta}{\lambda_{1}})=\tfrac{1}{2}\left(\frac{\delta+\sqrt{\delta}}{2}-\frac{2\delta}{\delta+\sqrt{\delta}}\right)$
$=\frac{\sqrt{\delta}}{2}\left(\frac{1+\sqrt{\delta}}{2}-\frac{2}{1+\sqrt{\delta}}\right)$
$=\frac{\sqrt{\delta}(-3+2\sqrt{\delta}+\delta)}{4(1+\sqrt{\delta})}$
$<0,$
from which it follows that $D_{1}<0$, and hence $D<0$. Thus the Pick matrix $P$ is not positive, and so, by Pick’s Theorem, the Pick interpolatioin problem $\lambda_{j}\mapsto z_{j},j=1,2$, is not solvable.
Proposition 5.24.
Let $\delta\in(0,1)$. The Szegő kernel $[\frac{1}{1-\bar{\mu}\lambda}]$ restricted to $R_{\delta}$ is not a DP Szegő kernel on $R_{\delta}$.
Proof.
Suppose the kernel $[\frac{1}{1-\bar{\mu}\lambda}]$ restricted to $R_{\delta}$ is a DP kernel. Then, for any distinct $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$, the localization of $[\frac{1}{1-\bar{\mu}\lambda}]$ to $\{\lambda_{1},\dots,\lambda_{n}\}$ belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$.
Consider the $2$ distinct points $\lambda_{1}=\frac{\delta+\sqrt{\delta}}{2}$, $\lambda_{2}=-\lambda_{1}$, note that, for $\delta\in(0,1)$, $\lambda_{1},\lambda_{2}\in R_{\delta}$. By Example (2.9), the function $\varphi(z)=\tfrac{1}{2}(z+\frac{\delta}{z})$ on $R_{\delta}$ satisfies $\|\varphi\|_{\mathrm{dp}}=1$. Therefore, for $\lambda_{i}$ and $z_{i}=\varphi(\lambda_{i})=\frac{1}{2}\left(\lambda_{i}+\frac{\delta}{\lambda_{i}}\right)$, $i=1,2$, the DP Pick problem
$\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,2,$ (5.25)
is solvable. In Example 5.22 we showed that, for $\delta\in(0,1)$, the corresponding Pick problem (5.25) is not solvable.
Since the problem (5.25) is a solvable DP Pick problem, by Theorem 5.2, for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$ (5.26)
By the assumption, the localization of $[\frac{1}{1-\bar{\mu}\lambda}]$ to $\{\lambda_{1},\lambda_{2}\}$ belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$. In particular, for the Pick problem
$\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,2,$ (5.27)
on $\mathbb{D}$, the Pick matrix
$\left[\frac{1-\bar{z}_{i}z_{j}}{1-\bar{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{2}\geq 0.$ (5.28)
Hence, by Pick’s theorem, the problem (5.27) is solvable by a Schur function $f$ on $\mathbb{D}$. This contradicts our example, and so $[\frac{1}{1-\bar{\mu}\lambda}]$ is not a DP Szegő kernel on $R_{\delta}$. ∎
6. Extremal DP Pick problems
We say that a DP Pick problem is extremally solvable if it is solvable and there does not exist $\varphi\in{\rm H}^{\infty}_{\rm dp}$ with $\|\varphi\|_{\rm dp}<1$ satisfying the equations
$\varphi(\lambda_{j})=z_{j}\qquad\text{for}\ j=1,\ldots,n.$ (6.1)
Remark 6.2.
A DP Pick problem that is not extremally solvable cannot have a unique solution. For suppose $\lambda_{j}\mapsto z_{j},j=1,\dots,n,$ is a solvable DP Pick problem that is not extremally solvable. That means that there is a function $\varphi:R_{\delta}\to\mathbb{C}$ such that $\varphi(\lambda_{j})=z_{j}$ for $j=1,\dots,n$ and $\|\varphi\|_{\mathrm{dp}}<1$. Consider the function $\psi(\lambda)=\varphi(\lambda)+\varepsilon\prod_{j=1}^{n}(\lambda-\lambda_{j})$, for $\lambda\in R_{\delta}$, for some positive $\varepsilon$. Then $\psi(\lambda_{j})=z_{j}$ for $j=1,\dots,n$ and
$\|\psi\|_{\mathrm{dp}}\leq\|\varphi\|_{\mathrm{dp}}+\varepsilon\|\prod_{j=1}^{n}(\lambda-\lambda_{j})\|_{\mathrm{dp}}<1$
for all small enough $\varepsilon$, and so there are infinitely many solutions to the interpolation problem $\lambda_{j}\mapsto z_{j}$ for $j=1,\dots,n$ having DP norm less than $1$.
Theorem 6.3.
Let $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$ be distinct and let $z_{1},\dots,z_{n}\in\mathbb{C}$. The following two statements are equivalent.
  1. (i)
    The DP Pick problem
    $\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,\ldots,n$
    is extremally solvable.
  2. (ii)
    For all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
    $[(1-\bar{z}_{i}z_{j})g_{ij}]_{i,j=1}^{n}\geq 0$ (6.4)
    and there exists $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
    $\operatorname{rank}\big{[}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\big{]}_{i,j=1}^{n}<n.$ (6.5)
Proof.
(i) $\implies$ (ii). Suppose that the DP Pick problem
$\lambda_{j}\mapsto z_{j}\ \text{for}\ j=1,\dots,n$
is extremally solvable. Since the problem is solvable, Theorem 5.2 implies that, for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$ (6.6)
Suppose, for a contradiction, that there is no $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ such that
$[(1-\bar{z}_{i}z_{j})g_{ij}]\ \text{is singular}.$ (6.7)
Let $F:\mathbb{R}\times\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)\to\mathbb{R}$ be defined by
[…]
$=\min_{J=1,\dots,n}\det\begin{bmatrix}(1-r^{2}\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{J}.$
By standard linear algebra, for any positive definite matrix $g=[g_{ij}]$, $F(r,g)>0$ if and only if $\begin{bmatrix}(1-r^{2}\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$. $F$ is continuous and, by supposition,
$[(1-\bar{z}_{i}z_{j})g_{ij}]>0$
for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$, which implies that $F(1,g)>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. Since, by Proposition 4.24, $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$ is compact, $F(1,\cdot)$ attains its minimum on $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, and so there exists $\kappa>0$ such that $F(1,g)\geq\kappa$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$.
By the continuity of $F$ and, again by the compactness of $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, the family of functions $\{F(\cdot,g):g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)\}$ is equicontinuous on $\mathbb{R}$. Hence there exists $\delta>0$ such that $F(r,g)>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$ and all $r\in(1,\delta)$. Choose some $r\in(1,\delta)$. Then $F(r,g)>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, and therefore $\begin{bmatrix}(1-r^{2}\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. It follows from Proposition 4.24 that $\begin{bmatrix}(1-r^{2}\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$. Hence, by Theorem 5.2, for the chosen $r\in(1,\delta)$, the DP Pick problem
$\lambda_{j}\mapsto rz_{j}\ \text{for}\ j=1,\dots,n$
is solvable, which is to say that there exists a function $\psi\in{\rm Hol}(R_{\delta})$ such that $\|\psi\|_{\mathrm{dp}}\leq 1$ and $\psi(\lambda_{j})=rz_{j}$ for $j=1,\dots,n$. Thus the function $\varphi\stackrel{{\scriptstyle\rm def}}{{=}}\psi/r$ satisfies $\varphi(\lambda_{j})=z_{j}$ for $j=1,\dots,n$ and $\|\varphi\|_{\mathrm{dp}}\leq 1/r<1$, contrary to hypothesis. Hence there is a $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ such that
$[(1-\bar{z}_{i}z_{j})g_{ij}]\ \text{is singular}.$ (6.8)
We have shown that statements (6.4) and (6.5) hold, and so have established (i) $\Rightarrow$ (ii) necessity in Theorem 6.3.
(ii) $\Rightarrow$ (i). Suppose that (ii) holds, and so, for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
$[(1-\bar{z}_{i}z_{j})g_{ij}]_{i,j=1}^{n}\geq 0.$ (6.9)
By Theorem 5.2, there exists $\varphi\in\mathcal{S}_{\mathrm{dp}}$ such that
$\varphi(\lambda_{j})=z_{j}\qquad\text{for}\ j=1,\ldots,n.$ (6.10)
Suppose (i) does not hold, which means that the problem is non-extremally solvable, and hence there exists $\varphi$ such that $\|\varphi\|_{\mathrm{dp}}=r<1$ and $\varphi$ satisfies $\varphi(\lambda_{j})=z_{j}$ for $j=1,\dots,n$. Thus for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
$[(r^{2}-\bar{z_{i}}z_{j})g_{ij}]\geq 0.$ (6.11)
By assumption (ii), there exists $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
$\operatorname{rank}\big{[}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\big{]}_{i,j=1}^{n}<n.$ (6.12)
and hence $[(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}]_{i,j=1}^{n}$ has a non-zero null vector $v$. Consider the relation
$(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}=(1-r^{2})\tilde{g}_{ij}+(r^{2}-\bar{z}_{i}z_{j})\tilde{g}_{ij}.$
Since $g_{ij}>0$, $(1-r^{2})\tilde{g}_{ij}>0$ and, by equation (6.11), $(r^{2}-\bar{z}_{i}z_{j})\tilde{g}_{ij}\geq 0$, which is a contradiction. ∎
Theorem 6.13.
Let $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$ be distinct and let $z_{1},\dots,z_{n}\in\mathbb{C}$. If the DP Pick problem
$\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,\ldots,n,$
is extremally solvable, then there exists a rational function $\varphi\in\mathcal{S}_{\mathrm{dp}}$ which satisfies the equations
$\varphi(\lambda_{j})=z_{j}\qquad\text{for}\ j=1,\ldots,n,$ (6.14)
and has a model $(\mathcal{M},u)$ as in equation (5.11), where $u:R_{\delta}\to\mathcal{M}$ is a holomorphic function and $\dim\mathcal{M}\leq 2n$. Furthermore, there exists $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$ such that
$1=\|\varphi\|_{\rm dp}=\|\varphi(T)\|.$
In particular,
$1-\varphi(T)^{*}\varphi(T)=u(T)^{*}\big{(}1-E(T)^{*}E(T)\big{)}\ u(T),$
where $\mathcal{M}$ can be written as $\mathcal{M}=\mathcal{M}_{1}\oplus\mathcal{M}_{2}$ and $E:R_{\delta}\to\mathcal{B}(\mathcal{M})$ is defined by the formula
$E(\lambda)=\begin{bmatrix}\lambda&0\\ 0&\frac{\delta}{\lambda}\end{bmatrix}\ \qquad\text{for}\ \lambda\in R_{\delta}$ (6.15)
with respect to this orthogonal decomposition of $\mathcal{M}$.
Proof.
Since the DP Pick problem $\lambda_{j}\mapsto z_{j},\ \text{for}\ j=1,\ldots,n,$ is solvable, by Theorem 5.9, there exists a rational function $\varphi\in\mathcal{S}_{\mathrm{dp}}$ such that $\varphi(\lambda_{j})=z_{j},\text{for}\ j=1,\ldots,n$.
Let us now prove the existence of an operator $T$ with the stated properties. By assumption, the DP-Pick problem $\lambda_{j}\mapsto z_{j},j=1,\dots,n$ is extremally solvable. By Theorem 6.3, there exists $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
$\operatorname{rank}\big{[}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\big{]}_{i,j=1}^{n}<n,$ (6.16)
so that $\begin{bmatrix}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\end{bmatrix}$ is singular, and therefore has a non-zero null vector $\xi=[\xi_{1},\dots,\xi_{n}]^{T}\in\mathbb{C}^{n}$, which is to say that
$\sum_{j=1}^{n}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\xi_{j}=0\ \text{for}\ i=1,\dots,n.$ (6.17)
Since $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$, $[\tilde{g}_{ij}]>0$, and so $[\tilde{g}_{ij}]$ has rank $n$. By Moore’s Theorem there exist an $n$-dimensional Hilbert space $\mathcal{H}$ and a basis $\tilde{e}_{1},\dots,\tilde{e}_{n}\in\mathcal{H}$ such that $\tilde{g}_{ij}=\langle\tilde{e}_{j},\tilde{e}_{i}\rangle$ for $i,j=1,\dots,n$.
Define an operator $T$ on $\mathcal{H}_{T}$ by $T\tilde{e}_{j}=\lambda_{j}\tilde{e}_{j}$ for $j=1,\dots,n$. Since $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$, by Proposition 4.10, $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$. Note that $\varphi(T)\tilde{e}_{j}=z_{j}\tilde{e}_{j}$, and so, if $x=\sum_{j=1}^{n}\xi_{j}\tilde{e}_{j}$, then
$\varphi(T)x=\sum_{j=1}^{n}z_{j}\xi_{j}\tilde{e}_{j}$
and
$\langle(1-\varphi(T)^{*}\varphi(T))x,x\rangle$ $=\sum_{i,j=1}^{n}(1-\bar{z}_{i}z_{j})\bar{\xi}_{i}\xi_{j}\langle\tilde{e}_{j},\tilde{e}_{i}\rangle$
$=\sum_{i,j=1}^{n}(1-\bar{z}_{i}z_{j})\bar{\xi}_{i}\xi_{j}\tilde{g}_{ij}$
$=\sum_{i=1}^{n}\bar{\xi}_{i}\sum_{j=1}^{n}(1-\bar{z}_{i}z_{j})\xi_{j}\tilde{g}_{ij}$
$=0.$ (6.18)
As $\xi\neq 0$, the complex numbers $\xi_{1},\dots,\xi_{n}$ are not all zero, and so, since $\tilde{e}_{1},\dots,\tilde{e}_{n}$ are linearly independent, $x=\sum_{j=1}^{n}\xi_{j}\tilde{e}_{j}\neq 0$. Since $\|\varphi\|_{\mathrm{dp}}\leq 1$ and $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$, we have $\|\varphi(T)\|\leq 1$, and so $1-\varphi(T)^{*}\varphi(T)\geq 0$. In conjunction with the inequality (6), this implies that $(1-\varphi(T)^{*}\varphi(T))x=0$, and hence $\|\varphi(T)x\|^{2}=\|x\|^{2}$. Since $x\neq 0$, $x$ is a maximizing vector for $\varphi(T)$ and $\|\varphi(T)\|=1$.
By Theorem 5.9, for the rational function $\varphi$ there exists a model $(\mathcal{M},u),$ where $u:R_{\delta}\to\mathcal{M}$ is holomorphic, so that
$1-\overline{\varphi(\mu)}\varphi(\lambda)=\Big{\langle}\,\big{(}1-E(\mu)^{*}E(\lambda)\big{)}\ u(\lambda)\,,u(\mu)\,\Big{\rangle}_{\mathcal{M}}\ \text{for}\ \lambda,\mu\in R_{\delta},$ (6.19)
where $\dim\mathcal{M}\leq 2n$. Since $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$, $T$ satisfies
$\sigma(T)=\{\lambda_{1},\ldots,\lambda_{n}\}\subset R_{\delta}.$
Thus, by the Riesz-Dunford functional calculus, $\varphi(T)$ is well defined and, by the hereditary functional calculus,
$1-\varphi(T)^{*}\varphi(T)=u(T)^{*}\big{(}1-E(T)^{*}E(T)\big{)}\ u(T).$

### v2

@@
 $R_{\delta}\stackrel{{\scriptstyle\rm def}}{{=}}\{z\in\mathbb{C}:\delta<|z|<1\},$
-where $10<\delta<$ 1.
+where $0<\delta<1$.
 The class of functions in question arises in the early work of R.
@@
 By analogy with the classical Schur class of holomorphic functions $\mathcal{S}$ with supremum norm at most $1$ on the disc $\mathbb{D}$, it is natural to consider the dp-Schur class $\mathcal{S}_{\mathrm{dp}}$ of holomorphic functions of dp-norm at most $1$ on $R_{\delta}$.
-Our central result is a generalization of the classical realization formula, for $\varphi\in\mathcal{S}$,
-$\varphi(z)=A+Bz(1-Dz)^{-1}C,\ \text{for}\ z\in\mathbb{D},$
-to functions from $\mathcal{S}_{\mathrm{dp}}$.
-A second result is a Pick interpolation theorem for functions in $\mathcal{S}_{\mathrm{dp}}$ that is analogous to Abrahamse’s Interpolation Theorem for bounded holomorphic functions on a multiply-connected domain.
+Our central result is a Pick interpolation theorem for functions in $\mathcal{S}_{\mathrm{dp}}$ that is analogous to Abrahamse’s Interpolation Theorem for bounded holomorphic functions on a multiply-connected domain.
 For a tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of distinct interpolation nodes in $R_{\delta}$, we introduce a special set $\mathcal{G}_{\mathrm{dp}}(\lambda)$ of positive definite $n\times n$ matrices, which we call DP Szegő kernels.
 The DP Pick problem $\lambda_{j}\mapsto z_{j},j=1,\dots,n$, is shown to be solvable if and only if,
-$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0\;\text{ for all}\;g\in\mathcal{G}_{\mathrm{dp}}(\lambda).$
+[…]
 We prove further that a solvable DP Pick problem has a solution which is a rational function with a finite-dimensional model, an intriguing result which opens up the possibility of a theory of extremal functions from $\mathcal{S}_{\mathrm{dp}}$ analogous to the theory of finite Blaschke products.
@@
 where
-$g_{\rho}(\lambda_{i},\lambda_{j})=\sum_{m=-\infty}^{\infty}\frac{(\bar{\lambda}_{i}\lambda_{j})^{m}}{\rho+\delta^{2m}},\quad\text{for}\ 1\leq i,j\leq n.$
+$g_{\rho}(\lambda_{i},\lambda_{j})=\sum_{m=-\infty}^{\infty}\frac{(\overline{\lambda}_{i}\lambda_{j})^{m}}{\rho+\delta^{2m}},\quad\text{for}\ 1\leq i,j\leq n.$
 Another natural variant of Pick’s problem arises if one replaces the supremum norm on ${\rm Hol}(\Omega)$ by a different norm.
 For example, consider the Dirichlet space $\mathcal{D}$ of holomorphic functions $f$ on $\mathbb{D}$ such that $f^{\prime}$ is square integrable with respect to area measure on $\mathbb{D}$, with pointwise operations and the norm
-[…]
+$\|f\|_{\mathcal{D}}^{2}=\sum_{n=0}^{\infty}(n+1)|\hat{f}(n)|^{2}=\|f\|_{H^{2}}^{2}+\int_{\mathbb{D}}|f^{\prime}(z)|^{2}dm(z),$
 where $m$ denotes area measure on the disc. The Dirichlet space is a Hilbert function space on $\mathbb{D}$ with reproducing kernel
-$k_{\mathcal{D}}(\lambda,\mu)=-\frac{1}{\bar{\mu}\lambda}\log(1-\bar{\mu}\lambda).$
+$k_{\mathcal{D}}(\lambda,\mu)=-\frac{1}{\overline{\mu}\lambda}\log(1-\overline{\mu}\lambda).$
 The Pick-type interpolation problem appropriate to this Hilbert function space is expressed in terms of its multiplier space $\mathcal{M}(\mathcal{D})$, which is defined to be the space of functions $\varphi$ on $\mathbb{D}$ such that $\varphi f\in\mathcal{D}$ for every $f\in\mathcal{D}$, with pointwise operations and the multiplier norm
@@
 $\|\varphi(X)\|\leq\left(2+\frac{1+\delta}{1-\delta}\right)\sup_{z\in R_{\delta}}\|\varphi(z)\|.$ (1.5)
-Hence, we see from equations (1.4) and (1.5) that $\|\varphi\|_{\mathrm{dp}}\leq\left(2+\frac{1+\delta}{1-\delta}\right)\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}$ for $\varphi\in{\rm Hol}(R_{\delta})$.
-On the other hand, trivially, $\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}\leq\|\varphi\|_{\mathrm{dp}}$, and so the dp and supremum norms on ${\rm Hol}(R_{\delta})$ are equivalent.
+Hence, we see from equations (1.4) and (1.5) that
+$\|\varphi\|_{\mathrm{dp}}\leq\left(2+\frac{1+\delta}{1-\delta}\right)\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}$
+for $\varphi\in{\rm Hol}(R_{\delta})$.
+On the other hand, see Remark 2.7, $\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}\leq\|\varphi\|_{\mathrm{dp}}$, and so the dp and supremum norms on ${\rm Hol}(R_{\delta})$ are equivalent.
 Thus,
@@
 In the scalar case a slightly stronger result than the inequality (1.5) had been obtained earlier by A.
-Shields [23, Proposition 23], with the smaller constant $2+\sqrt{\frac{1+\delta}{1-\delta}}$ on the right hand side.
+Shields [24, Proposition 23], with the smaller constant $2+\sqrt{\frac{1+\delta}{1-\delta}}$ on the right hand side.
 Shields asked whether the constant $2+\sqrt{\frac{1+\delta}{1-\delta}}$ could be replaced by a quantity that remains bounded as $\delta\to 1$.
@@
 Beckermann and M.
-Crouzeix [7] and subsequently a better constant was established by M.
-Crouzeix [11].
+Crouzeix [7] and subsequently the better constant $1+\sqrt{2}$ was established by M.
+Crouzeix and A.
+Greenbaum [11].
 Corresponding to the dp-Schur class there is a natural variant of the classical Pick interpolation problem, which we call the DP Pick problem: given $n$ distinct points $\lambda_{1},\ldots,\lambda_{n}$ in $R_{\delta}$ and $z_{1},\ldots,z_{n}\in\mathbb{C}$, does there exist a function $\varphi\in{\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ with $\|\varphi\|_{\mathrm{dp}}\leq 1$ such that
@@
 A DP Szegő kernel for the $n$-tuple $\lambda$ is a positive definite $n\times n$ matrix $g=[g_{ij}]$ such that
-$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]\geq 0\ \ \text{ and }\ \ \left[\left(1-\overline{\frac{\delta}{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]\geq 0.$ (1.9)
+[…]
 The set of all DP Szegő kernels for the $n$-tuple $\lambda$ will be denoted by $\mathcal{G}_{\mathrm{dp}}(\lambda)$.
-We observe that $\mathcal{G}_{\mathrm{dp}}(\lambda)$ consists of the gramians $[\langle e_{j},e_{i}\rangle]_{i,j=1}^{n}$ for all bases $e_{1},\dots,e_{n}$ of an $n$-dimensional Hilbert space $\mathcal{H}$ such that the operator $T$ on $\mathcal{H}$ defined by $Te_{j}=\lambda_{j}e_{j}$ for $j=1,\dots,n$ is a Douglas-Paulsen operator.
+We observe that $\mathcal{G}_{\mathrm{dp}}(\lambda)$ consists of the gramians $\langle e_{j},e_{i}\rangle]_{i,j=1}^{n}$ for all bases $e_{1},\dots,e_{n}$ of an $n$-dimensional Hilbert space $\mathcal{H}$ such that the operator $T$ on $\mathcal{H}$ defined by $Te_{j}=\lambda_{j}e_{j}$ for $j=1,\dots,n$ is a Douglas-Paulsen operator.
 This and related facts are described in Section 4.
@@
 if and only if, for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
-$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$ (1.11)
+$(1-\overline{z}_{i}z_{j})g_{ij}]\geq 0.$ (1.11)
 In Section 2 we compare the dp norm and the sup norm of a function in ${\rm Hol}(R_{\delta})$ and we point out a connection to the Crouzeix conjecture.
@@
 In this section we describe connections between the Banach algebra ${\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ and the Crouzeix conjecture.
-We will show in Proposition 2.11 that there is a large class of functions $\varphi\in{\rm Hol}(R_{\delta})$, such that
+We will prove in Proposition 2.11 that there is a large class of functions $\varphi\in{\rm Hol}(R_{\delta})$, such that
 $\|\varphi\|_{\mathrm{dp}}=\|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}.$
-A simple example of a function $f\in{\rm Hol}(R_{\delta})$ for which the last relation fails is $f(z)=z+\frac{\delta}{z}$, which satisfies
-$\|f\|_{\mathrm{dp}}=2\ \text{and}\ \|f\|_{{\rm H}^{\infty}(R_{\delta})}=1+\delta.$
+In Example 2.9 below we show that the last relation fails to hold for the function $\varphi\in{\rm Hol}(R_{\delta})$ defined by $\varphi(z)=z+\frac{\delta}{z}$ for $z\in R_{\delta}$.
+In fact $\varphi$ satisfies
+$\|\varphi\|_{\mathrm{dp}}=2\ \text{and}\ \|\varphi\|_{{\rm H}^{\infty}(R_{\delta})}=1+\delta.$
 By an elliptical domain we shall mean the domain in the complex plane bounded by an ellipse. As a standard elliptical domain we take the set
 $G_{\delta}\stackrel{{\scriptstyle\rm def}}{{=}}\{x+iy:x,y\in\mathbb{R},\frac{x^{2}}{(1+\delta)^{2}}+\frac{y^{2}}{(1-\delta)^{2}}<1\},$ (2.1)
-for some $\delta$ such that $10\leq\delta<$ 1.
+for some $\delta$ such that $0\leq\delta<1$.
 Note that any elliptical domain can be identified via an affine self-map of the plane with an elliptical domain of the form $G_{\delta}$ for some $\delta\in[0,1)$.
@@
 Delyon family, $\mathcal{F}_{\mathrm{bfd}}(C)$, corresponding to an open bounded convex set $C$ in $\mathbb{C}$ is the class of operators $T$ such that the closure of the numerical range of $T$, $\overline{W(T)}$, is contained in $C$.
-By [18, Theorem 1.2-1], the spectrum $\sigma(T)$ of an operator $T$ is contained in $\overline{W(T)}$, and so, by the Riesz-Dunford functional calculus, $\varphi(T)$ is defined for all $\varphi\in{\rm Hol}(C)$ and $T\in\mathcal{F}_{\mathrm{bfd}}(C)$.
+By [19, Theorem 1.2-1], the spectrum $\sigma(T)$ of an operator $T$ is contained in $\overline{W(T)}$, and so, by the Riesz-Dunford functional calculus, $\varphi(T)$ is defined for all $\varphi\in{\rm Hol}(C)$ and $T\in\mathcal{F}_{\mathrm{bfd}}(C)$.
 Therefore, we may consider the calcular norm [^3] A calcular norm on a function space is a norm that is defined with the aid of the functional calculus.
@@
 $K_{\mathrm{bfd}}=\sup\{K(\mathcal{F}_{\mathrm{bfd}}(C)):\ C\ \text{is a bounded convex set in }\ \mathbb{C}\}.$
-In [9], Crouzeix proved $K_{\mathrm{bfd}}\leq 12$ and conjectured that $K_{\mathrm{bfd}}=2$. Subsequently Crouzeix and Palencia [10] proved that $K_{\mathrm{bfd}}\leq 1+\sqrt{2}$.
+In [9], Crouzeix proved $K_{\mathrm{bfd}}\leq 12$ and conjectured that $K_{\mathrm{bfd}}=2$.
+Subsequently Crouzeix and Palencia [10] proved that $K_{\mathrm{bfd}}\leq 1+\sqrt{2}$.
+Still more recently Crouzeix and Kressner [12] showed that $W(T)$ is a complete $(1+\sqrt{2})$-spectral set for $T$.
 Let $\pi:R_{\delta}\to G_{\delta}$ be defined by $\pi(z)=z+\frac{\delta}{z}$, $z\in R_{\delta}$.
@@
 $\|f\|_{{\rm H}^{\infty}(R_{\delta})}=\sup_{z\in R_{\delta}}\left|z+\frac{\delta}{z}\right|=1+\delta.$
-Note that $\varphi(z)=z$ has bfd-norm equal to $2$ and sup norm on $G_{\delta}$ equal to $1+\delta$. Hence the Crouzeix universal constant $K_{\mathrm{bfd}}\geq 2$. ∎
+Note that $\varphi(z)=z$ has bfd-norm equal to 22 and sup norm on $G_{\delta}$ equal to $1+\delta$. Hence the Crouzeix universal constant $K_{\mathrm{bfd}}\geq 2$. ∎
 Remark 2.10.
-The inequality $K_{\mathrm{bfd}}\geq 2$ was proved by G. Tsikalas [22]. He considered the domain $A_{R}=\{1/R<|z|<R\}$ where $R>1,$ and the functions $g_{n}$ in ${\rm Hol}(A_{R})$ defined by
-$g_{n}(z)=\frac{1}{R^{n}}\left(\frac{1}{z^{n}}+z^{n}\right),\ n=1,2,\dots.$
-He showed that
-$\limsup_{n\to\infty}\frac{\|g_{n}\|_{\mathrm{dp}}}{\|g_{n}\|_{{\rm H}^{\infty}(A_{R})}}\geq 2,$
-from which it follows that $K_{\mathrm{bfd}}\geq 2$.
+In [23] G.
+Tsikalas proved a result about the annulus as a $K$-spectral set.
+We restate his result in the notation of this paper as follows.
+Let $K(\delta)$ denote the smallest constant such that $R_{\delta}$ is a $K(\delta)$-spectral set for any bounded linear operator $T\in\mathcal{F}_{\rm dp}(\delta)$.
+He used the functions $g_{n}$ in ${\rm Hol}(R_{\delta})$ defined by
+$g_{n}(z)=\frac{\delta^{n}}{z^{n}}+z^{n},\quad\text{for}\ n=1,2,\dots,$
+to show that $K(\delta)\geq 2$, for all $\delta\in(0,1)$.
 Proposition 2.11.
@@
 $\|\varphi\|_{\mathrm{dp}}$ $=\sup_{X\in\mathcal{F}_{\mathrm{dp}}(\delta)}\|\varphi(X)\|$
-$\leq\sup_{\|X\|\leq 1}\|\varphi(X)\|$ by the definition of $\text{by the definition of }\mathcal{F}_{\mathrm{dp}}(\delta)$
-$=\|\varphi\|_{{\rm H}^{\infty}(\mathbb{D})}.$ by von Neumann’s inequality (2.13)
+$\leq\sup_{\|X\|\leq 1}\|\varphi(X)\|\$$\text{by the definition of }\mathcal{F}_{\mathrm{dp}}(\delta)$
+$=\|\varphi\|_{{\rm H}^{\infty}(\mathbb{D})}.\$by von Neumann’s inequality (2.13)
 By the Maximum principle, for $\varphi\in{\rm Hol}(\mathbb{D})$,
@@
 3. Models and realizations of holomorphic functions on $R_{\delta}$
-In this section we review some known results on the function theory of holomorphic functions in the dp-norm on an annulas.
+In this section we review some known results on the function theory of holomorphic functions in the dp-norm on an annulus.
 The models and realizations of holomorphic functions $\varphi:R_{\delta}\to\mathbb{C}$ such that $\|\varphi\|_{\mathrm{dp}}\leq 1$ are presented in [6, Theorem 9.46].
@@
 There exists a $\mathrm{dp}$-model $(\mathcal{N},v)$ of $\varphi$ with parameter $\delta$, in the sense that there are Hilbert spaces $\mathcal{N}^{+},\mathcal{N}^{-}$ and an ordered pair $v=(v^{+},v^{-})$ of holomorphic functions, where $v^{+}:R_{\delta}\to\mathcal{N}^{+}$ and $v^{-}:R_{\delta}\to\mathcal{N}^{-}$ satisfy, for all $z,w\in R_{\delta}$,
-$1-\overline{\varphi(w)}\varphi(z)=(1-\bar{w}z)\langle v^{+}(z),v^{+}(w)\rangle_{\mathcal{N}^{+}}+(\bar{w}z-\delta^{2})\langle v^{-}(z),v^{-}(w)\rangle_{\mathcal{N}^{-}}.$
+$1-\overline{\varphi(w)}\varphi(z)=(1-\overline{w}z)\langle v^{+}(z),v^{+}(w)\rangle_{\mathcal{N}^{+}}+(\overline{w}z-\delta^{2})\langle v^{-}(z),v^{-}(w)\rangle_{\mathcal{N}^{-}}.$
 Definition 3.2.
-A positive semi-definite function on a set $X$ is a function $A:X\times X\to\mathbb{C}$ such that, for any positive integer $n$ and any points $x_{1},\dots,x_{n}\in X$, the $n\times n$ matrix $[A(x_{j},x_{i})]_{i,j=1}^{n}$ is positive semi-definite.
+A positive semi-definite function on a set $X$ is a function $A:X\times X\to\mathbb{C}$ such that, for any positive integer $n$ and any points $x_{1},\dots,x_{n}\in X$, the $n\times n$ matrix $A(x_{j},x_{i})]_{i,j=1}^{n}$ is positive semi-definite.
 We shall write
-$[A(x,y)]\geq 0,\ \text{for all}\ x,y\in X,$
+$A(x,y)]\geq 0,\ \text{for all}\ x,y\in X,$
 to mean that $A$ is a positive semi-definite function on $X$.
@@
 $\varphi\in\mathcal{S}_{\mathrm{dp}}$ if and only if there exist a pair of positive semi-definite functions $A$ and $B$ on $R_{\delta}$ such that
-$1-\overline{\varphi(\mu)}\varphi(\lambda)=(1-\bar{\mu}\lambda)A(\lambda,\mu)+(1-\overline{\frac{\delta}{\mu}}\frac{\delta}{\lambda})B(\lambda,\mu)$ (3.4)
+$1-\overline{\varphi(\mu)}\varphi(\lambda)=(1-\overline{\mu}\lambda)A(\lambda,\mu)+(1-\frac{\delta}{\overline{\mu}}\frac{\delta}{\lambda})B(\lambda,\mu)$ (3.4)
 for all $\lambda,\mu\in R_{\delta}$.
@@
 where $\mathcal{M}$ satisfying (3.5) is chosen to have minimal dimension. [^5] Equivalently, $\{u(\lambda)\,:\,\lambda\in\Omega\}$ spans $\mathcal{M}$.
-Definition 3.8.
-If $\varphi\in{\rm Hol}(R_{\delta})$ we say $\varphi$ is of finite type if $\varphi\in\mathcal{S}_{\mathrm{dp}}$ and, in addition, $\varphi$ can be represented as in equation (3.4) where $A$ and $B$ have finite rank.
+The following theorem is stated as [6, Theorem 9.54]. For the convenience of the reader we shall give a full proof here.
 Theorem 3.8.
@@
 As explained in Theorem 3.1, it means that there exist Hilbert spaces $\mathcal{N}^{+}$ and $\mathcal{N}^{-}$ and maps $v^{+}:R_{\delta}\to\mathcal{N}^{+},v^{-}:R_{\delta}\to\mathcal{N}^{-}$ such that, for all $\lambda,\mu\in R_{\delta}$,
-$1-\overline{\varphi(\mu)}\varphi(\lambda)=(1-\bar{\mu}\lambda)\langle v^{+}(\lambda),v^{+}(\mu)\rangle_{\mathcal{N}^{+}}+(\bar{\mu}\lambda-\delta^{2})\langle v^{-}(\lambda),v^{-}(\mu)\rangle_{\mathcal{N}^{-}}.$
+$1-\overline{\varphi(\mu)}\varphi(\lambda)=(1-\overline{\mu}\lambda)\langle v^{+}(\lambda),v^{+}(\mu)\rangle_{\mathcal{N}^{+}}+(\overline{\mu}\lambda-\delta^{2})\langle v^{-}(\lambda),v^{-}(\mu)\rangle_{\mathcal{N}^{-}}.$
 Reshuffle this relation to
@@
 for all $\lambda\in R_{\delta}$.
-Since both $\mathcal{X}$ and $\mathcal{Y}$ are subspaces of $\mathbb{C}\oplus\mathcal{N}$, we may extend $L$ (possibly after enlarging the space $\mathcal{N}$) to an isometry $L:\mathcal{N}\to\mathcal{N}$ (see the discussion in [6, Remark 2.31] for this step).
+Since both $\mathcal{X}$ and $\mathcal{Y}$ are subspaces of $\mathbb{C}\oplus\mathcal{N}$, we may extend $L$ (possibly after enlarging the space $\mathcal{N}$) to a unitary operator $L:\mathcal{N}\to\mathcal{N}$ (see the discussion in [6, Remark 2.31] for this step).
 Write $L$ as a block operator matrix
@@
 $\gamma+DE(\lambda)u(\lambda)$ $=u(\lambda),$ (3.15)
-where $E(\lambda)$ is given by equation (3.6).
-Since $\|D\|\leq 1$ and $\|E(\lambda)\|=\max\{|\lambda|,\frac{\delta}{|\lambda|}\}<1$ for all $\lambda\in R_{\delta}$, it follows that $1-DE(\lambda)$ is invertible for $\lambda\in R_{\delta}$, and hence
+where $E(\lambda)$ is given by equation (3.6). Since $\|D\|\leq 1$ and
+$\|E(\lambda)\|=\max{\left\{|\lambda|,\frac{\delta}{|\lambda|}\right\}}<1\ \text{ for all}\ \lambda\in R_{\delta},$
+it follows that $1-DE(\lambda)$ is invertible for $\lambda\in R_{\delta}$, and hence
 $u(\lambda)$ $=(1-DE(\lambda))^{-1}\gamma,$
@@
 for all $\lambda\in R_{\delta}$, which is the desired realization formula (3.10).
-Conversely, suppose that $a,\beta,\gamma.D$ are such that $L$ given by equation (3.14) is an isometry on $\mathbb{C}\oplus\mathcal{N}$ and that $\varphi$ is the function on $R_{\delta}$ defined by equation (3.11).
+Conversely, suppose that $a,\beta,\gamma,D$ are such that $L$ given by equation (3.13) is a unitary operator on $\mathbb{C}\oplus\mathcal{N}$ and that $\varphi$ is the function on $R_{\delta}$ defined by equation (3.10).
 Since $1-DE(\lambda)$ is invertible for all $\lambda\in R_{\delta}$ we may define a mapping $u:R_{\delta}\to\mathcal{N}$ by equation (3.11).
@@
 if and only if there exist a pair of $n\times n$ positive semi-definite matrices $A=[a_{ij}]$ and $B=[b_{ij}]$ such that
-$1-\overline{z_{i}}z_{j}=(1-\bar{\lambda}_{i}\lambda_{j})a_{ij}+(1-\frac{\delta^{2}}{\bar{\lambda}_{i}\lambda_{j}})b_{ij}$
+$1-\overline{z_{i}}z_{j}=(1-\overline{\lambda}_{i}\lambda_{j})a_{ij}+(1-\frac{\delta^{2}}{\overline{\lambda}_{i}\lambda_{j}})b_{ij}$
 for $i,j=1,\dots,n$.
@@
 In this section we follow Abrahamse’s idea of using families of kernels to solve Pick interpolation problems.
+To this end we shall introduce several objects that depend on an $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$ of distinct points in $R_{\delta}$.
+First we consider the set $\mathcal{F}_{\rm dp}(\delta,\lambda)$ of operators on $n$-dimensional Hilbert space with spectrum $\{\lambda_{1},\dots,\lambda_{n}\}$ which belong to the Douglas-Paulsen family $\mathcal{F}_{\rm dp}(\delta)$.
+Secondly we define DP Szegő kernels for the $n$-tuple $\lambda$.
+We establish a close connection between these two objects in Propositions 4.9 and 4.10.
+Thereby, in Section 5 we shall establish a theorem analogous to Theorem 1.2, Abrahamse’s Theorem.
 Definition 4.1.
 We say that a kernel $k:R_{\delta}\times R_{\delta}$ is a DP Szegő kernel on $R_{\delta}$ if
-$[(1-\bar{\mu}\lambda)k(\lambda,\mu)]\geq 0\ \ \text{ and }\ \ [(1-\frac{\delta}{\overline{\mu}}\frac{\delta}{\lambda})k(\lambda,\mu)]\geq 0,\qquad\text{for all}\;\lambda,\mu\in R_{\delta}.$ (4.2)
+[…]
 We let
@@
 Let $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ be an $n$-tuple of distinct points in $R_{\delta}$.
-We denote by $\mathcal{F}_{\rm dp}(\delta,\lambda)$ the set of operators $T$ in the Douglas-Paulsen family $\mathcal{F}_{\mathrm{dp}}(\delta)$ corresponding to the annulus $R_{\delta}$ that act on an $n$-dimensional Hilbert space $\mathcal{H}_{T}$ and satisfy
+We denote by $\mathcal{F}_{\rm dp}(\delta,\lambda)$ the family of operators $T$ in the Douglas-Paulsen family $\mathcal{F}_{\mathrm{dp}}(\delta)$ corresponding to the annulus $R_{\delta}$ that act on an $n$-dimensional Hilbert space $\mathcal{H}_{T}$ and satisfy
 $\sigma(T)=\{\lambda_{1},\ldots,\lambda_{n}\}.$
@@
 Then, we shall prove in Proposition 4.9 that $g=[g_{ij}]$ is a positive definite $n\times n$ matrix such that
-$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]\geq 0\ \ \text{ and }\ \ \left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]\geq 0.$ (4.6)
+[…]
 Definition 4.7.
@@
 We define $\mathcal{G}_{\mathrm{dp}}(\lambda)$ to be the set of positive definite $n\times n$ matrices $g=[g_{ij}]$ such that
-$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]\geq 0\ \ \text{ and }\ \ \left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]\geq 0.$ (4.8)
+[…]
 We call $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ a DP-Szegő kernel for the $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$.
@@
 By assumption $T$ is a Douglas-Paulsen operator with parameter $\delta$ that acts on an $n$-dimensional Hilbert space $\mathcal{H}_{T}$, $T$ has $n$ linearly independent eigenvectors $e_{1},\dots,e_{n}$ corresponding to the eigenvalues $\lambda_{1},\dots,\lambda_{n}$ respectively and $g_{ij}=\langle e_{j},e_{i}\rangle$ for $i,j=1,\dots,n$.
-By the definition of the Douglas-Paulsen class, $\|T\|\leq 1$ and $\|\delta/T\|\leq 1$, so that, for any vector $x=\sum_{j=1}^{n}x_{j}e_{j}$, we have
-00 0 $\leq\|x\|^{2}-\|Tx\|^{2}$
+By the definition of the Douglas-Paulsen class, $\|T\|\leq 1$ and $\|\delta T^{-1}\|\leq 1$, so that, for any vector $x=\sum_{j=1}^{n}x_{j}e_{j}$, we have
+$0$ $\leq\|x\|^{2}-\|Tx\|^{2}$
 $=\langle\sum_{j=1}^{n}x_{j}e_{j},\sum_{i=1}^{n}x_{i}e_{i}\rangle-\langle\sum_{j=1}^{n}x_{j}\lambda_{j}e_{j},\sum_{i=1}^{n}\lambda_{i}x_{i}e_{i}\rangle$
@@
 Thus
-$\big{[}(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}\big{]}_{i,j=1}^{n}\geq 0.$
-Likewise, the relation $\|\frac{\delta}{T}x\|\leq\|x\|$ holds for any vector $x=\sum_{j=1}^{n}x_{j}e_{j}\in\mathcal{H}_{T}$. Therefore, we have
-00 0 $\leq\|x\|^{2}-\left\|\frac{\delta}{T}x\right\|^{2}$
+$\big[(1-\overline{\lambda}_{i}\lambda_{j})g_{ij}\big]_{i,j=1}^{n}\geq 0.\$
+Likewise, the relation $\|\delta T^{-1}x\|\leq\|x\|$ holds for any vector $x=\sum_{j=1}^{n}x_{j}e_{j}\in\mathcal{H}_{T}$. Therefore, we have
+$0$ $\leq\|x\|^{2}-\left\|\delta T^{-1}x\right\|^{2}$
 $=\langle\sum_{j=1}^{n}x_{j}e_{j},\sum_{i=1}^{n}x_{i}e_{i}\rangle-\langle\sum_{j=1}^{n}\frac{\delta}{\lambda_{j}}x_{j}e_{j},\sum_{i=1}^{n}\frac{\delta}{\lambda_{i}}x_{i}e_{i}\rangle$
@@
 $\|Tx\|^{2}$ $=\big\langle\,\sum_{j=1}^{n}\xi_{j}\lambda_{j}e_{j},\sum_{i=1}^{n}\xi_{i}\lambda_{i}e_{i}\,\big\rangle$
-$=\sum_{j,i=1}^{n}\xi_{j}\lambda_{j}\bar{\xi}_{i}\bar{\lambda}_{i}\langle\,e_{j},e_{i}\,\rangle$
-$=\sum_{j,i=1}^{n}\bar{\lambda}_{i}\lambda_{j}\xi_{j}\bar{\xi}_{i}\,g_{ij}.$ (4.12)
+$=\sum_{j,i=1}^{n}\xi_{j}\lambda_{j}\overline{\xi}_{i}\overline{\lambda}_{i}\langle\,e_{j},e_{i}\,\rangle$
+$=\sum_{j,i=1}^{n}\overline{\lambda}_{i}\lambda_{j}\xi_{j}\overline{\xi}_{i}\,g_{ij}.$ (4.12)
 Hence
-$\|x\|^{2}-\|Tx\|^{2}=\sum_{j,i=1}^{n}(1-\bar{\lambda}_{i}\lambda_{j})\,g_{ij}\xi_{j}\bar{\xi}_{i}.$
+$\|x\|^{2}-\|Tx\|^{2}=\sum_{j,i=1}^{n}(1-\overline{\lambda}_{i}\lambda_{j})\,g_{ij}\xi_{j}\overline{\xi}_{i}.$
 By hypothesis,
-$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]\geq 0,$
+$(1-\overline{\lambda}_{i}\lambda_{j})g_{ij}]\geq 0,$
 and so $\|x\|^{2}-\|Tx\|^{2}\geq 0$. Thus $\|T\|\leq 1$. Similarly, using the hypothesis
@@
 Let $\lambda_{1},\ldots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$ and $z_{1},\ldots,z_{n}\in\mathbb{C}$.
-If the DP Pick Problem 3.17 is solvable, then, for any positive definite $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
-$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$ (4.15)
+If the DP Pick Problem 3.16 is solvable, then, for any positive definite $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
+$(1-\overline{z}_{i}z_{j})g_{ij}]\geq 0.$ (4.14)
 Proof.
-By assumption, the DP Pick Problem 3.17 is solvable, that is, there exist a function $\varphi\in{\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ with $\|\varphi\|_{\mathrm{dp}}\leq 1$ and satisfying
+By assumption, the DP Pick Problem 3.16 is solvable, that is, there exist a function $\varphi\in{\rm H}^{\infty}_{\mathrm{dp}}(R_{\delta})$ with $\|\varphi\|_{\mathrm{dp}}\leq 1$ and satisfying
 $\varphi(\lambda_{j})=z_{j},\qquad j=1,\ldots,n.$ (4.15)
@@
 Therefore, by equation (4), the condition $\|\varphi(T)\|\leq 1$ translates into
-$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$
+$(1-\overline{z}_{i}z_{j})g_{ij}]\geq 0.\qed$
 Definition 4.18.
-Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. We say that a DP-Szegő kernel $[g_{ij}]\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ is normalized if $g_{ii}=1$ for $i=1,\dots,n$.
+Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$. We say that a DP-Szegő kernel $g_{ij}]\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ is normalized if $g_{ii}=1$ for $i=1,\dots,n$.
 Let $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$ denote the set of normalized DP-Szegő kernels for the $n$-tuple $(\lambda_{1},\dots,\lambda_{n})$.
@@
 Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$.
-Every DP-Szegő kernel $[g_{ij}]$ from $\mathcal{G}_{\mathrm{dp}}(\lambda)$ is diagonally congruent to a normalized DP-Szegő kernel.
+Every DP-Szegő kernel $g_{ij}]$ from $\mathcal{G}_{\mathrm{dp}}(\lambda)$ is diagonally congruent to a normalized DP-Szegő kernel.
 Proof.
-For any matrix $[g_{ij}]\in\mathcal{G}_{\mathrm{dp}}(\lambda)$, we can define a positive definite matrix $[h_{ij}]$ by
+For any matrix $g_{ij}]\in\mathcal{G}_{\mathrm{dp}}(\lambda)$, we can define a positive definite matrix $h_{ij}]$ by
 $h_{ii}=1\ \text{for}\ i=1,\dots,n\ \text{and}\ h_{ij}=c_{i}^{-1}g_{ij}c_{j}^{-1}\ \text{if}\ i\neq j$ (4.20)
@@
 $\begin{bmatrix}h_{ij}\end{bmatrix}_{i,j=1}^{n}=C^{*}\begin{bmatrix}g_{ij}\end{bmatrix}_{i,j=1}^{n}C\ \text{where}\ C=\mathrm{diag}\{1/c_{1},\dots,1/c_{n}\}.$ (4.22)
-On conjugating the inequalities (4.8) by the matrix $C$ we find that $[h_{ij}]$ belongs to $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. ∎
+On conjugating the inequalities (4.8) by the matrix $C$ we find that $h_{ij}]$ belongs to $\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$. ∎
 Proposition 4.23.
@@
 Moreover, for fixed target data $z_{1},\dots,z_{n}$,
-$\begin{bmatrix}(1-\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}\geq 0\ \text{for all}\ g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ (4.25)
+$\begin{bmatrix}(1-\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}\geq 0\ \text{for all}\ g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ (4.24)
 if and only if
-$\begin{bmatrix}(1-\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}\geq 0\ \text{for all}\ g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda).$ (4.26)
+$\begin{bmatrix}(1-\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}\geq 0\ \text{for all}\ g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda).$ (4.25)
 Proof.
@@
 Then $h\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, and so, by assumption,
-$[(1-\bar{z}_{i}z_{j})h_{ij}]\geq 0$ (4.31)
+$(1-\overline{z}_{i}z_{j})h_{ij}]\geq 0$ (4.30)
 Conjugate this matrix inequality by $\mathrm{diag}\{c_{1},\dots,c_{n}\}$ to obtain the relation (4.24). Thus the relation (4.25) implies the relation (4.24). ∎
-Definition 4.32.
-Let $\lambda_{1},\dots,\lambda_{n}$ be $n$ distinct points in $R_{\delta}$.
-Let $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ be a DP Szegő kernel for the $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$.
-The matrices
-$[(1-\bar{\lambda}_{i}\lambda_{j})g_{ij}]_{i,j=1}^{n}\ \ \text{ and }\ \ \left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)g_{ij}\right]_{i,j=1}^{n}$ (4.33)
-are said to be the defect matrices of $g$.
 Say that a DP-Szegő kernel $g$ on $R_{\delta}$ is reducible if there exist DP-Szegő kernels $h$ and $k$ on $R_{\delta}$ such that $g=h+k$ and neither $h$ nor $k$ is diagonally congruent to $g$.
@@
 Clearly, if DP Pick data $\lambda_{j}\mapsto z_{j},j=1,\dots,n$, are such that
-$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0\ \text{and}\ [1-\frac{\delta^{2}}{\overline{z_{i}}z_{j}}g_{ij}]\geq 0$
+$(1-\overline{z}_{i}z_{j})g_{ij}]\geq 0\ \text{and}\ [1-\frac{\delta^{2}}{\overline{z_{i}}z_{j}}g_{ij}]\geq 0$
 for all irreducible DP Szegő kernels $g$ then the same inequality holds for all DP Szegő kernels, and consequently the DP pick interpolation problem is solvable.
 Since the class of irreducible DP Szegő kernels is likely to be much smaller than the class of all DP Szegő kernels, it would be valuable to identify the irreducible DP Szegő kernels on $R_{\delta}$.
+Problem 4.31.
+Find an effective description of the irreducible DP Szegő kernels on $R_{\delta}$.
 5. The DP Pick problem and DP-Szegő kernels
+In this section we shall prove our main theorem, which is a solvability criterion for DP Pick problems in terms of DP-Szegő kernels.
+We also present some examples which illustrate the relationship between the Pick and DP Pick interpolation problems.
+The following notation and terminology will be needed in the proofs.
 Definition 5.1.
@@
 if and only if, for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
-$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$ (5.3)
+$(1-\overline{z}_{i}z_{j})g_{ij}]\geq 0.$ (5.3)
 Proof.
@@
 To prove $\Leftarrow$, suppose that
-$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0$ (5.4)
+$(1-\overline{z}_{i}z_{j})g_{ij}]\geq 0$ (5.4)
 for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$.
 By Theorem 3.18, to show that the DP Pick Problem (3.16) is solvable it suffices to prove that there exist a pair of $n\times n$ positive semi-definite matrices $A=[a_{ij}]$ and $B=[b_{ij}]$ such that
-$1-\overline{z_{i}}z_{j}=(1-\bar{\lambda}_{i}\lambda_{j})a_{ij}+(\bar{\lambda}_{i}\lambda_{j}-\delta^{2})b_{ij}$
+$1-\overline{z_{i}}z_{j}=(1-\overline{\lambda_{i}}\lambda_{j})a_{ij}+(\overline{\lambda}_{i}\lambda_{j}-\delta^{2})b_{ij}$
 for all $i,j=1,\dots,n$. Let $H_{n}$ be the real linear space of Hermitian matrices in $\mathbb{C}^{n\times n}$, and let
-${\mathcal{C}}=\left\{[(1-\bar{\lambda}_{i}\lambda_{j})a_{ij}]_{i,j=1}^{n}++\left[\left(1-\frac{\delta^{2}}{\bar{\lambda}_{i}\lambda_{j}}\right)b_{ij}\right]_{i,j=1}^{n}:\,[a_{ij}]_{i,j=1}^{n}\geq 0\ \text{and}\ [b_{ij}]_{i,j=1}^{n}\geq 0\right\}\subset H_{n}.$ (5.5)
+${\mathcal{C}}=\left\{[(1-\overline{\lambda_{i}}\lambda_{j})a_{ij}]_{i,j=1}^{n}+\left[\left(1-\frac{\delta^{2}}{\overline{\lambda_{i}}\lambda_{j}}\right)b_{ij}\right]_{i,j=1}^{n}:[a_{ij}]_{i,j=1}^{n}\geq 0\ \text{and}\ [b_{ij}]_{i,j=1}^{n}\geq 0\right\}.$ (5.5)
 The subset $\mathcal{C}$} is a closed convex cone in $H_{n}$.
-Note that every $n\times n$ positive semi-definite matrix $[a_{ij}]_{i,j=1}^{n}$ belongs to ${\mathcal{C}}$.
-By the positivity of Szegő kernel $\left[\frac{1}{1-\bar{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{n}$, the $n\times n$ matrix of the form
-$\left[\frac{a_{ij}}{1-\bar{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{n}$
+Note that every $n\times n$ positive semi-definite matrix $a_{ij}]_{i,j=1}^{n}$ belongs to $\mathcal{C}$}.
+By the positivity of Szegő kernel $\left[\frac{1}{1-\overline{\lambda_{i}}\lambda_{j}}\right]_{i,j=1}^{n}$, the $n\times n$ matrix of the form
+$\left[\frac{a_{ij}}{1-\overline{\lambda_{i}}\lambda_{j}}\right]_{i,j=1}^{n}$
 is also positive semi-definite.
-In the definition of ${\mathcal{C}}$ (5.5) we may replace $[a_{ij}]_{i,j=1}^{n}$ by $\left[\frac{a_{ij}}{1-\bar{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{n}$ and $[b_{ij}]_{i,j=1}^{n}$ by the zero matrix, to deduce that $[a_{ij}]_{i,j=1}^{n}$ belongs to ${\mathcal{C}}$.
-By the Hahn-Banach theorem, to show that $[1-\bar{z}_{i}z_{j}]_{i,j=1}^{n}$ belongs to ${\mathcal{C}}$ it suffices to prove that, for every real linear functional ${\mathcal{L}}$ on $H_{n}$, ${\mathcal{L}}\geq 0$ on ${\mathcal{C}}$ implies ${\mathcal{L}}([1-\bar{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0$.
-Extend ${\mathcal{L}}$ to a complex linear functional $\tilde{\mathcal{L}}$ on $\mathbb{C}^{n\times n}$ by
-~$\tilde{\mathcal{L}}(X+{\mathrm{i}}\,Y)={\mathcal{L}}(X)+{\mathrm{i}}\,{\mathcal{L}}(Y)$
+In the definition of $\mathcal{C}$} (5.5) we may replace $a_{ij}]_{i,j=1}^{n}$ by $\left[\frac{a_{ij}}{1-\overline{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{n}$ and $b_{ij}]_{i,j=1}^{n}$ by the zero matrix, to deduce that $a_{ij}]_{i,j=1}^{n}$ belongs to $\mathcal{C}$}.
+By the Hahn-Banach theorem, to show that $1-\overline{z_{i}}z_{j}]_{i,j=1}^{n}$ belongs to $\mathcal{C}$} it suffices to prove that, for every real linear functional $\mathcal{L}$} on $H_{n}$, ${\mathcal{L}}\geq 0$ on $\mathcal{C}$} implies ${\mathcal{L}}([1-\overline{z_{i}}z_{j}]_{i,j=1}^{n})\geq 0$.
+Extend $\mathcal{L}$} to a complex linear functional $\widetilde{\mathcal{L}}$ on $\mathbb{C}^{n\times n}$ by
+$\widetilde{\mathcal{L}}(X+{\mathrm{i}}\,Y)={\mathcal{L}}(X)+{\mathrm{i}}\,{\mathcal{L}}(Y)$
 for $X,Y\in H_{n}$. Now define a pre-inner product $\langle\cdot,\cdot\rangle_{L}$ on $\mathbb{C}^{n}$ by
-$\langle\,c,d\,\rangle_{L}=\tilde{\mathcal{L}}(c\otimes d)$
+$\langle\,c,d\,\rangle_{L}=\widetilde{\mathcal{L}}(c\otimes d)$
 for $c,d\in\mathbb{C}^{n}$. Here $c\otimes d\in\mathbb{C}^{n\times n}$, defined by
@@
 Note that, for any $c\in\mathbb{C}^{n}$,
-$\langle\,c,c\,\rangle_{L}=\tilde{\mathcal{L}}(c\otimes c)={\mathcal{L}}(c\otimes c)\geq 0.$
+$\langle\,c,c\,\rangle_{L}=\widetilde{\mathcal{L}}(c\otimes c)={\mathcal{L}}(c\otimes c)\geq 0.$
 Let
@@
 $Te_{j}=\lambda_{j}e_{j},\ j=1,\dots,n.$ (5.6)
+Let us construct an operator $\widetilde{T}$ on $\mathbb{C}^{n}/{\mathcal{N}}$ such that $\|\widetilde{T}\|\leq 1$ and $\|\delta\widetilde{T}^{-1}\|\leq 1$.
 For $x=\sum_{j=1}^{n}\xi_{j}e_{j}\in\mathbb{C}^{n}$, we have
-$\langle\,x,x\,\rangle_{L}-\langle\,Tx,Tx\,\rangle_{L}$ =~$=\tilde{\mathcal{L}}(x\otimes x)-\tilde{\mathcal{L}}(Tx\otimes Tx)$
-=~$=\tilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}e_{i}\right)-\tilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}\lambda_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}\lambda_{i}e_{i}\right)$
-=~$=\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\bar{\xi}_{i}e_{j}\,\otimes e_{i}\right)-\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\lambda_{j}\bar{\xi}_{i}\bar{\lambda}_{i}e_{j}\,\otimes e_{i}\right)$
-=~$=\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}(1-\bar{\lambda}_{i}\lambda_{j})\bar{\xi}_{i}\xi_{j}e_{j}\,\otimes e_{i}\right)$
-=~$=\tilde{\mathcal{L}}\left[(1-\bar{\lambda}_{i}\lambda_{j})\bar{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}$
-$={\mathcal{L}}\left[(1-\bar{\lambda}_{i}\lambda_{j})\bar{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}\geq 0\;\;\;\text{since}\;\;\mathcal{L}\geq 0\ \text{on}\ {\mathcal{C}}.$ (5.7)
+$\langle\,x,x\,\rangle_{L}-\langle\,Tx,Tx\,\rangle_{L}$ $=\widetilde{\mathcal{L}}(x\otimes x)-\widetilde{\mathcal{L}}(Tx\otimes Tx)$
+$=\widetilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}e_{i}\right)-\widetilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}\lambda_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}\lambda_{i}e_{i}\right)$
+$=\widetilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\overline{\xi}_{i}e_{j}\,\otimes e_{i}\right)-\widetilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\lambda_{j}\overline{\xi}_{i}\overline{\lambda}_{i}e_{j}\,\otimes e_{i}\right)$
+$=\widetilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}(1-\overline{\lambda}_{i}\lambda_{j})\overline{\xi}_{i}\xi_{j}e_{j}\,\otimes e_{i}\right)$
+$=\widetilde{\mathcal{L}}\left[(1-\overline{\lambda}_{i}\lambda_{j})\overline{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}$
+$={\mathcal{L}}\left[(1-\overline{\lambda}_{i}\lambda_{j})\overline{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}\geq 0\;\;\;\text{since}\;\;\mathcal{L}\geq 0\ \text{on}\ {\mathcal{C}}.$ (5.7)
 Thus
-$\langle\,x,x\,\rangle_{L}-\langle\,Tx,Tx\,\rangle_{L}\geq 0,$
-and so $x\in{\mathcal{N}}$ implies that $Tx\in{\mathcal{N}}$. Hence $T$ induces an operator $\tilde{T}$ on $\mathbb{C}^{n}/{\mathcal{N}}$ by
-$\tilde{T}(x+{\mathcal{N}})=Tx+{\mathcal{N}},$
-and $\|\tilde{T}(x+{\mathcal{N}})\|^{2}\leq\|x+{\mathcal{N}}\|^{2}$ for all $(x+{\mathcal{N}})\in\mathbb{C}^{n}/{\mathcal{N}}$.
-Similarly, let us prove that $\|\delta\tilde{T}^{-1}\|\leq 1$. For $x=\sum_{j=1}^{n}\xi_{j}e_{j}\in\mathbb{C}^{n}$, we have
-$\langle\,x,x\,\rangle_{L}-\langle\,\delta T^{-1}x,\delta T^{-1}x\,\rangle_{L}$ =~$=\tilde{\mathcal{L}}(x\otimes x)-\tilde{\mathcal{L}}(\delta T^{-1}x\otimes\delta T^{-1}x)$
-=~$=\tilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}e_{i}\right)-\tilde{\mathcal{L}}\left(\sum_{j=1}^{n}\xi_{j}\frac{\delta}{\lambda_{j}}e_{j}\,\otimes\sum_{i=1}^{n}\xi_{i}\frac{\delta}{\lambda_{i}}e_{i}\right)$
-=~$=\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\bar{\xi}_{i}e_{j}\,\otimes e_{i}\right)-\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\xi_{j}\frac{\delta}{\lambda_{j}}\bar{\xi}_{i}\frac{\delta}{\overline{\lambda_{i}}}e_{j}\,\otimes e_{i}\right)$
-=~$=\tilde{\mathcal{L}}\left(\sum_{j,i=1}^{n}\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)\bar{\xi}_{i}\xi_{j}e_{j}\,\otimes e_{i}\right)$
-=~$=\tilde{\mathcal{L}}\left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)\bar{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}$
-$={\mathcal{L}}\left[\left(1-\frac{\delta}{\overline{\lambda_{i}}}\frac{\delta}{\lambda_{j}}\right)\bar{\xi}_{i}\xi_{j}\right]_{j,i=1}^{n}\geq 0\;\;\;\text{since}\;\;\mathcal{L}\geq 0\ \text{on}\ {\mathcal{C}}.$ (5.8)
-Thus
-$\langle\,x,x\,\rangle_{L}-\langle\,\delta T^{-1}x,\delta T^{-1}x\,\rangle_{L}\geq 0,$
-and so $x\in{\mathcal{N}}$ implies that $\delta T^{-1}x\in{\mathcal{N}}$.
-Therefore, $T$ is a Douglas-Paulsen operator.
-Since the eigenvalues of $T$, $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$, $\sigma(T)\subseteq R_{\delta}$, the operator $T\in\mathcal{F}_{\mathrm{dp}}(\delta)$.
-Therefore, by Proposition 4.9, $[\langle\,e_{j},e_{i}\,\rangle_{L}]_{i,j=1}^{n}$ is a DP Szegő kernel for the $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})$.
-By supposition,
-$[(1-\bar{z}_{i}z_{j})\langle\,e_{j},e_{i}\,\rangle_{L}]_{i,j=1}^{n}\geq 0.$
-Choose a polynomial $p$ such that $p(\lambda_{i})=z_{i}$, $i=1,\dots,n$.
-Then ${}p(\tilde{T})e_{i}=z_{i}e_{i}$, $i=1,\dots,n$, and so $\|p(\tilde{T})\|\leq 1$.
-Choose $c=\begin{bmatrix}1\\ \dots\\ 1\end{bmatrix}\in\mathbb{C}^{n}$.
-Then
-$\langle\,(1-p(\tilde{T})^{*}p(\tilde{T}))c,c\,\rangle_{L}\geq 0,$
+$\langle\,x,x\,\rangle_{L}-\langle\,Tx,Tx\,\rangle_{L}\geq 0,$ (5.8)
+and so $x\in{\mathcal{N}}$ implies that $Tx\in{\mathcal{N}}$. Hence $T$ induces an operator $\widetilde{T}$ on $\mathbb{C}^{n}/{\mathcal{N}}$ by
+$\widetilde{T}(x+{\mathcal{N}})=Tx+{\mathcal{N}},$
+and $\|\widetilde{T}(x+{\mathcal{N}})\|^{2}\leq\|x+{\mathcal{N}}\|^{2}$ for all $(x+{\mathcal{N}})\in\mathbb{C}^{n}/{\mathcal{N}}$, which implies that
+$\|\widetilde{T}\|\leq 1.$ (5.9)
+Notice that $\sigma(T)=\{\lambda_{1},\dots,\lambda_{n}\}\subset R_{\delta}$ and so $T$ is invertible. Moreover
+$\delta T^{-1}e_{j}=\frac{\delta}{\lambda_{j}}e_{j}\ \text{for}\ j=1,\dots,n,$
+and so, in the chain of equations leading to equation (5), we may replace $T$ by $\delta T^{-1}$ and $\lambda_{j}$ by $\frac{\delta}{\lambda_{j}}$ to deduce that, for $x=\sum_{j=1}^{n}\xi_{j}e_{j}\in\mathbb{C}^{n}$,
+$\langle x,x\rangle_{L}-\langle\delta T^{-1}x,\delta T^{-1}x\rangle_{L}=\mathcal{L}\left[\left(1-\tfrac{\delta}{\overline{\lambda_{i}}}\tfrac{\delta}{\lambda_{j}}\right)\overline{\xi_{i}}\xi_{j}\right]_{j,i=1}^{n}.$
+Clearly $\left[\left(1-\tfrac{\delta}{\overline{\lambda_{i}}}\tfrac{\delta}{\lambda_{j}}\right)\overline{\xi_{i}}\xi_{j}\right]_{j,i=1}^{n}\in\mathcal{C}$ (take $a_{ij}=0,b_{ij}=\overline{\xi_{i}}\xi_{j}$ in the defining expression (5.5)), and so, since $\mathcal{L}\geq 0$ on $\mathcal{C}$, we have
+$\langle x,x\rangle_{L}-\langle\delta T^{-1}x,\delta T^{-1}x\rangle_{L}\geq 0.$ (5.10)
+Thus $x\in{\mathcal{N}}$ implies that $\delta T^{-1}x\in{\mathcal{N}}$, and therefore $\delta T^{-1}$ induces an operator $\widetilde{(\delta T^{-1})}$ on $\mathbb{C}^{n}/\mathcal{N}$ by
+$\widetilde{(\delta T^{-1})}(x+\mathcal{N})=\delta T^{-1}x+\mathcal{N},$
+and in the light of inequality (5.10),
+$\|\widetilde{(\delta T^{-1})}\|\leq 1.$ (5.11)
+We have, for any $x\in\mathbb{C}^{n}$,
+$\widetilde{T}\widetilde{(\delta T^{-1})}(x+\mathcal{N})=\widetilde{T}(\delta T^{-1}x+\mathcal{N})=T\delta T^{-1}x+\mathcal{N}=\delta(x+\mathcal{N}),$
+and so $\widetilde{(\delta T^{-1})}=\delta(\widetilde{T})^{-1}$. Hence, by the inequality (5.11),
+$\|\delta(\widetilde{T})^{-1}\|=\|\widetilde{(\delta T^{-1})}\|\leq 1.$
+Therefore, $\tilde{T}$ is a Douglas-Paulsen operator.
+Since the eigenvalues of $T$, which are $\lambda_{1},\dots,\lambda_{n}$, belong to $R_{\delta}$, $\sigma(T)\subseteq R_{\delta}$, and so the operator $T$ belongs to $\mathcal{F}_{\mathrm{dp}}(\delta,\lambda)$.
+Therefore, by Proposition 4.9, $\langle\,e_{j},e_{i}\,\rangle_{L}]_{i,j=1}^{n}$ belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$.
+Let $g_{ij}=\langle\,e_{j},e_{i}\,\rangle_{L}$ for $i,j=1,\dots,n$. By supposition (5.4),
+$(1-\overline{z}_{i}z_{j})\langle\,e_{j},e_{i}\,\rangle_{L}]_{i,j=1}^{n}\geq 0.$
+Choose a polynomial $p$ such that $p(\lambda_{i})=z_{i}$, $i=1,\dots,n$. Then $p(\widetilde{T})e_{i}=z_{i}e_{i}$, $i=1,\dots,n$. Observe that
+$\left[\langle(1-p(\tilde{T})^{*}p(\tilde{T}))e_{j},e_{i}\rangle\right]$ $=\left[\langle e_{j},e_{i}\rangle-\langle p(\tilde{T})e_{j},p(\tilde{T})e_{i}\rangle\right]$
+$=\left[\langle e_{j},e_{i}\rangle-\langle z_{j}e_{j},z_{i}e_{i}\rangle\right]$
+$=\left[(1-\overline{z}_{i}z_{j})g_{ij}\right]\geq 0.$
+Therefore, $\|p(\widetilde{T})\|\leq 1$. Choose $c=\begin{bmatrix}1\\ \dots\\ 1\end{bmatrix}\in\mathbb{C}^{n}$. Then
+$\langle\,(1-p(\widetilde{T})^{*}p(\widetilde{T}))c,c\,\rangle_{L}\geq 0,$
 that is,
-${\mathcal{L}}([1-\bar{z}_{i}z_{j}]_{i,j=1}^{n}*cc^{*})\geq 0,\ \text{and so }\ {\mathcal{L}}([1-\bar{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0,$
+${\mathcal{L}}([1-\overline{z}_{i}z_{j}]_{i,j=1}^{n}*cc^{*})\geq 0,\ \text{and so }\ {\mathcal{L}}([1-\overline{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0,$
 where * denotes the Schur product of matrices.
 Thus, for every real linear functional $\mathcal{L}$} on $H_{n}$ such that ${\mathcal{L}}\geq 0$ on $\mathcal{C}$} we have
-${\mathcal{L}}([1-\bar{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0.$
-Hence $[1-\bar{z}_{i}z_{j}]_{i,j=1}^{n}$ belongs to ${\mathcal{C}}$. ∎
+${\mathcal{L}}([1-\overline{z}_{i}z_{j}]_{i,j=1}^{n})\geq 0.$
+Hence $1-\overline{z}_{i}z_{j}]_{i,j=1}^{n}$ belongs to $\mathcal{C}$}. ∎
+We show in the next theorem that, as in the classical Pick theorem, if a DP Pick problem is solvable then it is solvable by a rational function in $\mathcal{S}_{\mathrm{dp}}$.
 Theorem 5.12.
@@
 is a solvable DP-Pick problem. By Theorem 3.18, there exist positive semi-definite $n\times n$ matrices $a=\begin{bmatrix}a_{ij}\end{bmatrix}$ and $b=\begin{bmatrix}b_{ij}\end{bmatrix}$ such that
-$1-\bar{z}_{i}z_{j}=(1-\bar{\lambda}_{i}\lambda_{j})a_{ij}+(1-\frac{\delta^{2}}{\overline{\lambda_{i}}\lambda_{j}})b_{ij}\quad\text{for}\ i,j=1,\dots,n.$ (5.13)
+$1-\overline{z}_{i}z_{j}=(1-\overline{\lambda}_{i}\lambda_{j})a_{ij}+(1-\frac{\delta^{2}}{\overline{\lambda_{i}}\lambda_{j}})b_{ij}\quad\text{for}\ i,j=1,\dots,n.$ (5.16)
 Let the ranks of the matrices $a,b$ be $r_{1},r_{2}$ respectively, so that $r_{1}\leq n,r_{2}\leq n$.
@@
 Substituting these relations into the equations (5.16) and re-arranging, we obtain the relations
-$1+\langle\lambda_{j}x_{j},\lambda_{i}x_{i}\rangle_{\mathbb{C}^{r_{1}}}+\langle\frac{\delta}{\lambda_{j}}y_{j},\frac{\delta}{\lambda_{i}}y_{i}\rangle_{\mathbb{C}^{r_{2}}}=\bar{z}_{i}z_{j}+\langle x_{j},x_{i}\rangle_{\mathbb{C}^{r_{1}}}+\langle y_{j},y_{i}\rangle_{\mathbb{C}^{r_{2}}}\ \text{for}\ i,j=1,\dots,n.$
+$1+\langle\lambda_{j}x_{j},\lambda_{i}x_{i}\rangle_{\mathbb{C}^{r_{1}}}+\langle\frac{\delta}{\lambda_{j}}y_{j},\frac{\delta}{\lambda_{i}}y_{i}\rangle_{\mathbb{C}^{r_{2}}}=\overline{z}_{i}z_{j}+\langle x_{j},x_{i}\rangle_{\mathbb{C}^{r_{1}}}+\langle y_{j},y_{i}\rangle_{\mathbb{C}^{r_{2}}}\ \text{for}\ i,j=1,\dots,n.$
 These equations can in turn be expressed by saying that the families of vectors
@@
 and so the restriction of $\varphi$ to $R_{\delta}$ is in $\mathcal{S}_{\mathrm{dp}}$, which is to say that the corresponding DP Pick problem is solvable. ∎
+As the dp norm and sup norm are different, the converse statement to Remark 5.21 is false, as one would expect. The following two examples provide concrete instances of this fact.
 Example 5.23.
@@
 $\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,2,$
-we have $3$ cases:
+solvability depends on the value of $\delta$. There are $3$ cases:
   1. (i)
@@
 $\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,2,$
-is not solvable, while, for $\delta=\tfrac{1}{4}$, the Pick interpolation problem is uniquely solvable, and one sees by inspection that the unique solution is the function $f(\lambda)=\lambda$.
+is not solvable, while, for $\delta=\tfrac{1}{4}$, the Pick interpolation problem is uniquely solvable, and one sees by inspection that the unique solution is the function $f(\lambda)=\lambda$. ∎
 Example 5.25.
-A solvable DP-Pick problem on $R_{\delta}$ which is not solvable as a Pick problem on $\mathbb{D}$.
+Another solvable DP Pick data-set which is not a solvable Pick data-set on $\mathbb{D}$.
 Let $\delta\in(0,1)$ and let $\lambda_{1}=\frac{\delta+\sqrt{\delta}}{2}$, $\lambda_{2}=-\lambda_{1}$.
-We have $10<\delta<\frac{\delta+\sqrt{\delta}}{2}<\sqrt{\delta}<$ 1, so that $\lambda_{1},\lambda_{2}\in R_{\delta}$.
+We have $0<\delta<\frac{\delta+\sqrt{\delta}}{2}<\sqrt{\delta}<1$, so that $\lambda_{1},\lambda_{2}\in R_{\delta}$.
 Recall from Example 2.9 that the function $\varphi(z)=\tfrac{1}{2}(z+\frac{\delta}{z})$ on $R_{\delta}$ satisfies $\|\varphi\|_{\mathrm{dp}}=1$.
@@
 from which it follows that $D_{1}<0$, and hence $D<0$.
-Thus the Pick matrix $P$ is not positive, and so, by Pick’s Theorem, the Pick interpolatioin problem $\lambda_{j}\mapsto z_{j},j=1,2$, is not solvable.
+Thus the Pick matrix $P$ is not positive, and so, by Pick’s Theorem, the Pick interpolation problem $\lambda_{j}\mapsto z_{j},j=1,2$, is not solvable.
+Since the Pick interpolation problem on $\mathbb{D}$ and the DP Pick problem on $R_{\delta}$ are so closely related, it is natural to ask whether the Szegő kernel on $\mathbb{D}$, when retricted to $R_{\delta}$, is a DP-Szegő kernel.
+We can use Example 5.25 to answer this question in the negative.
 Proposition 5.27.
-Let $\delta\in(0,1)$. The Szegő kernel $[\frac{1}{1-\bar{\mu}\lambda}]$ restricted to $R_{\delta}$ is not a DP Szegő kernel on $R_{\delta}$.
+Let $\delta\in(0,1)$. The Szegő kernel $\frac{1}{1-\overline{\mu}\lambda}]$ restricted to $R_{\delta}$ is not a DP Szegő kernel on $R_{\delta}$.
 Proof.
-Suppose the kernel $[\frac{1}{1-\bar{\mu}\lambda}]$ restricted to $R_{\delta}$ is a DP kernel.
-Then, for any distinct $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$, the localization of $[\frac{1}{1-\bar{\mu}\lambda}]$ to $\{\lambda_{1},\dots,\lambda_{n}\}$ belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$.
+Suppose the kernel $\frac{1}{1-\overline{\mu}\lambda}]$ restricted to $R_{\delta}$ is a DP kernel.
+Then, for any distinct $\lambda_{1},\dots,\lambda_{n}\in R_{\delta}$, the localization of $\frac{1}{1-\overline{\mu}\lambda}]$ to $\{\lambda_{1},\dots,\lambda_{n}\}$ belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$.
 Consider the $2$ distinct points $\lambda_{1}=\frac{\delta+\sqrt{\delta}}{2}$, $\lambda_{2}=-\lambda_{1}$, note that, for $\delta\in(0,1)$, $\lambda_{1},\lambda_{2}\in R_{\delta}$.
-By Example (2.9), the function $\varphi(z)=\tfrac{1}{2}(z+\frac{\delta}{z})$ on $R_{\delta}$ satisfies $\|\varphi\|_{\mathrm{dp}}=1$.
+By Example 2.9, the function $\varphi(z)=\tfrac{1}{2}(z+\frac{\delta}{z})$ on $R_{\delta}$ satisfies $\|\varphi\|_{\mathrm{dp}}=1$.
 Therefore, for $\lambda_{i}$ and $z_{i}=\varphi(\lambda_{i})=\frac{1}{2}\left(\lambda_{i}+\frac{\delta}{\lambda_{i}}\right)$, $i=1,2$, the DP Pick problem
@@
 Since the problem (5.28) is a solvable DP Pick problem, by Theorem 5.2, for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
-$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$ (5.26)
-By the assumption, the localization of $[\frac{1}{1-\bar{\mu}\lambda}]$ to $\{\lambda_{1},\lambda_{2}\}$ belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$. In particular, for the Pick problem
+$(1-\overline{z}_{i}z_{j})g_{ij}]\geq 0.$ (5.29)
+By the assumption, the localization of $\frac{1}{1-\overline{\mu}\lambda}]$ to $\{\lambda_{1},\lambda_{2}\}$ belongs to $\mathcal{G}_{\mathrm{dp}}(\lambda)$. In particular, for the Pick problem
 $\lambda_{j}\mapsto z_{j}\quad\text{for}\ j=1,2,$ (5.30)
 on $\mathbb{D}$, the Pick matrix
-$\left[\frac{1-\bar{z}_{i}z_{j}}{1-\bar{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{2}\geq 0.$ (5.28)
+$\left[\frac{1-\overline{z}_{i}z_{j}}{1-\overline{\lambda}_{i}\lambda_{j}}\right]_{i,j=1}^{2}\geq 0.$ (5.31)
 Hence, by Pick’s theorem, the problem (5.30) is solvable by a Schur function $f$ on $\mathbb{D}$.
-This contradicts our example, and so $[\frac{1}{1-\bar{\mu}\lambda}]$ is not a DP Szegő kernel on $R_{\delta}$.
+This contradicts our example, and so $\frac{1}{1-\overline{\mu}\lambda}]$ is not a DP Szegő kernel on $R_{\delta}$.
 ∎
 6. Extremal DP Pick problems
+In this section we study DP Pick interpolation problems that are “only just” solvable.
 We say that a DP Pick problem is extremally solvable if it is solvable and there does not exist $\varphi\in{\rm H}^{\infty}_{\rm dp}$ with $\|\varphi\|_{\rm dp}<1$ satisfying the equations
@@
 for all small enough $\varepsilon$, and so there are infinitely many solutions to the interpolation problem $\lambda_{j}\mapsto z_{j}$ for $j=1,\dots,n$ having DP norm less than $1$.
+Next we give necessary and sufficient conditions for a DP Pick problem to be extremally solvable.
 Theorem 6.3.
@@
     For all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
-    $[(1-\bar{z}_{i}z_{j})g_{ij}]_{i,j=1}^{n}\geq 0$ (6.4)
-    and there exists $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
-    $\operatorname{rank}\big{[}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\big{]}_{i,j=1}^{n}<n.$ (6.5)
+    $(1-\overline{z}_{i}z_{j})g_{ij}]_{i,j=1}^{n}\geq 0$ (6.4)
+    and there exists $\widetilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
+    $\operatorname{rank}\big[(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}\big]_{i,j=1}^{n}<n.$ (6.5)
 Proof.
@@
 is extremally solvable. Since the problem is solvable, Theorem 5.2 implies that, for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
-$[(1-\bar{z}_{i}z_{j})g_{ij}]\geq 0.$ (6.6)
+$(1-\overline{z}_{i}z_{j})g_{ij}]\geq 0.$ (6.6)
 Suppose, for a contradiction, that there is no $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ such that
-$[(1-\bar{z}_{i}z_{j})g_{ij}]\ \text{is singular}.$ (6.7)
+[…]
 Let $F:\mathbb{R}\times\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)\to\mathbb{R}$ be defined by
-[…]
-$=\min_{J=1,\dots,n}\det\begin{bmatrix}(1-r^{2}\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{J}.$
-By standard linear algebra, for any positive definite matrix $g=[g_{ij}]$, $F(r,g)>0$ if and only if $\begin{bmatrix}(1-r^{2}\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$.
+$F(r,[g_{ij}])$ $=\text{the minimum of the leading principal minors of}\begin{bmatrix}(1-r^{2}\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}$
+$=\min_{J=1,\dots,n}\det\begin{bmatrix}(1-r^{2}\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{J}.$
+By standard linear algebra, for any positive definite matrix $g=[g_{ij}]$, $F(r,g)>0$ if and only if $\begin{bmatrix}(1-r^{2}\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$.
 $F$ is continuous and, by supposition,
-$[(1-\bar{z}_{i}z_{j})g_{ij}]>0$
+$(1-\overline{z}_{i}z_{j})g_{ij}]>0$
 for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$, which implies that $F(1,g)>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$.
@@
 Choose some $r\in(1,\delta)$.
-Then $F(r,g)>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, and therefore $\begin{bmatrix}(1-r^{2}\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$.
-It follows from Proposition 4.24 that $\begin{bmatrix}(1-r^{2}\bar{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$.
+Then $F(r,g)>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$, and therefore $\begin{bmatrix}(1-r^{2}\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}^{\mathrm{norm}}(\lambda)$.
+It follows from Proposition 4.23 that $\begin{bmatrix}(1-r^{2}\overline{z}_{i}z_{j})g_{ij}\end{bmatrix}_{i,j=1}^{n}>0$ for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$.
 Hence, by Theorem 5.2, for the chosen $r\in(1,\delta)$, the DP Pick problem
@@
 Hence there is a $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$ such that
-$[(1-\bar{z}_{i}z_{j})g_{ij}]\ \text{is singular}.$ (6.8)
+[…]
 We have shown that statements (6.4) and (6.5) hold, and so have established (i) $\Rightarrow$ (ii) necessity in Theorem 6.3.
 (ii) $\Rightarrow$ (i). Suppose that (ii) holds, and so, for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
-$[(1-\bar{z}_{i}z_{j})g_{ij}]_{i,j=1}^{n}\geq 0.$ (6.9)
+$(1-\overline{z}_{i}z_{j})g_{ij}]_{i,j=1}^{n}\geq 0.$ (6.9)
 By Theorem 5.2, there exists $\varphi\in\mathcal{S}_{\mathrm{dp}}$ such that
@@
 Thus for all $g\in\mathcal{G}_{\mathrm{dp}}(\lambda)$,
-$[(r^{2}-\bar{z_{i}}z_{j})g_{ij}]\geq 0.$ (6.11)
-By assumption (ii), there exists $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
-$\operatorname{rank}\big{[}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\big{]}_{i,j=1}^{n}<n.$ (6.12)
-and hence $[(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}]_{i,j=1}^{n}$ has a non-zero null vector $v$. Consider the relation
-$(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}=(1-r^{2})\tilde{g}_{ij}+(r^{2}-\bar{z}_{i}z_{j})\tilde{g}_{ij}.$
-Since $g_{ij}>0$, $(1-r^{2})\tilde{g}_{ij}>0$ and, by equation (6.11), $(r^{2}-\bar{z}_{i}z_{j})\tilde{g}_{ij}\geq 0$, which is a contradiction. ∎
+$(r^{2}-\overline{z_{i}}z_{j})g_{ij}]\geq 0.$ (6.11)
+By assumption (ii), there exists $\widetilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
+$\operatorname{rank}\big[(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}\big]_{i,j=1}^{n}<n.$ (6.12)
+and hence $(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}]_{i,j=1}^{n}$ has a non-zero null vector $v$. Consider the relation
+$(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}=(1-r^{2})\widetilde{g}_{ij}+(r^{2}-\overline{z}_{i}z_{j})\widetilde{g}_{ij}.$
+Since $\widetilde{g}_{ij}>0$, $(1-r^{2})\widetilde{g}_{ij}>0$ and, by equation (6.11), $(r^{2}-\overline{z}_{i}z_{j})\widetilde{g}_{ij}\geq 0$, which is a contradiction. ∎
+By Theorem 5.12, if a DP Pick problem is solvable, then there exists a rational solution $\varphi\in\mathcal{S}_{\mathrm{dp}}$.
+In the next theorem we show that if, further, the problem is extremally solvable then there exists $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$, acting on an $n$-dimensional Hilbert space, such that $\|\varphi(T)\|=\|\varphi\|_{\rm dp}=1.$
 Theorem 6.13.
@@
 By assumption, the DP-Pick problem $\lambda_{j}\mapsto z_{j},j=1,\dots,n$ is extremally solvable.
-By Theorem 6.3, there exists $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
-$\operatorname{rank}\big{[}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\big{]}_{i,j=1}^{n}<n,$ (6.16)
-so that $\begin{bmatrix}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\end{bmatrix}$ is singular, and therefore has a non-zero null vector $\xi=[\xi_{1},\dots,\xi_{n}]^{T}\in\mathbb{C}^{n}$, which is to say that
-$\sum_{j=1}^{n}(1-\bar{z}_{i}z_{j})\tilde{g}_{ij}\xi_{j}=0\ \text{for}\ i=1,\dots,n.$ (6.17)
-Since $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$, $[\tilde{g}_{ij}]>0$, and so $[\tilde{g}_{ij}]$ has rank $n$.
-By Moore’s Theorem there exist an $n$-dimensional Hilbert space $\mathcal{H}$ and a basis $\tilde{e}_{1},\dots,\tilde{e}_{n}\in\mathcal{H}$ such that $\tilde{g}_{ij}=\langle\tilde{e}_{j},\tilde{e}_{i}\rangle$ for $i,j=1,\dots,n$.
-Define an operator $T$ on $\mathcal{H}_{T}$ by $T\tilde{e}_{j}=\lambda_{j}\tilde{e}_{j}$ for $j=1,\dots,n$.
-Since $\tilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$, by Proposition 4.10, $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$.
-Note that $\varphi(T)\tilde{e}_{j}=z_{j}\tilde{e}_{j}$, and so, if $x=\sum_{j=1}^{n}\xi_{j}\tilde{e}_{j}$, then
-$\varphi(T)x=\sum_{j=1}^{n}z_{j}\xi_{j}\tilde{e}_{j}$
+By Theorem 6.3, there exists $\widetilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$ such that
+$\operatorname{rank}\big[(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}\big]_{i,j=1}^{n}<n,$ (6.16)
+so that $\begin{bmatrix}(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}\end{bmatrix}$ is singular, and therefore has a non-zero null vector $\xi=[\xi_{1},\dots,\xi_{n}]^{T}\in\mathbb{C}^{n}$, which is to say that
+$\sum_{j=1}^{n}(1-\overline{z}_{i}z_{j})\widetilde{g}_{ij}\xi_{j}=0\ \text{for}\ i=1,\dots,n.$ (6.17)
+Since $\widetilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$, $\widetilde{g}_{ij}]>0$, and so $\widetilde{g}_{ij}]$ has rank $n$.
+By Moore’s theorem’s Theorem there exist an $n$-dimensional Hilbert space $\mathcal{H}$ and a basis $\widetilde{e}_{1},\dots,\widetilde{e}_{n}\in\mathcal{H}$ such that $\widetilde{g}_{ij}=\langle\widetilde{e}_{j},\widetilde{e}_{i}\rangle$ for $i,j=1,\dots,n$.
+Define an operator $T$ on $\mathcal{H}$ by $T\widetilde{e}_{j}=\lambda_{j}\widetilde{e}_{j}$ for $j=1,\dots,n$.
+Since $\widetilde{g}\in\mathcal{G}_{\rm dp}(\lambda)$, by Proposition 4.10, $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$.
+Note that $\varphi(T)\widetilde{e}_{j}=z_{j}\widetilde{e}_{j}$, and so, if $x=\sum_{j=1}^{n}\xi_{j}\widetilde{e}_{j}$, then
+$\varphi(T)x=\sum_{j=1}^{n}z_{j}\xi_{j}\widetilde{e}_{j}$
 and
-$\langle(1-\varphi(T)^{*}\varphi(T))x,x\rangle$ $=\sum_{i,j=1}^{n}(1-\bar{z}_{i}z_{j})\bar{\xi}_{i}\xi_{j}\langle\tilde{e}_{j},\tilde{e}_{i}\rangle$
-$=\sum_{i,j=1}^{n}(1-\bar{z}_{i}z_{j})\bar{\xi}_{i}\xi_{j}\tilde{g}_{ij}$
-$=\sum_{i=1}^{n}\bar{\xi}_{i}\sum_{j=1}^{n}(1-\bar{z}_{i}z_{j})\xi_{j}\tilde{g}_{ij}$
+$\langle(1-\varphi(T)^{*}\varphi(T))x,x\rangle$ $=\sum_{i,j=1}^{n}(1-\overline{z}_{i}z_{j})\overline{\xi}_{i}\xi_{j}\langle\widetilde{e}_{j},\widetilde{e}_{i}\rangle$
+$=\sum_{i,j=1}^{n}(1-\overline{z}_{i}z_{j})\overline{\xi}_{i}\xi_{j}\widetilde{g}_{ij}$
+$=\sum_{i=1}^{n}\overline{\xi}_{i}\sum_{j=1}^{n}(1-\overline{z}_{i}z_{j})\xi_{j}\widetilde{g}_{ij}$
 $=0.$ (6.18)
-As $\xi\neq 0$, the complex numbers $\xi_{1},\dots,\xi_{n}$ are not all zero, and so, since $\tilde{e}_{1},\dots,\tilde{e}_{n}$ are linearly independent, $x=\sum_{j=1}^{n}\xi_{j}\tilde{e}_{j}\neq 0$.
+As $\xi\neq 0$, the complex numbers $\xi_{1},\dots,\xi_{n}$ are not all zero, and so, since $\widetilde{e}_{1},\dots,\widetilde{e}_{n}$ are linearly independent, $x=\sum_{j=1}^{n}\xi_{j}\widetilde{e}_{j}\neq 0$.
 Since $\|\varphi\|_{\mathrm{dp}}\leq 1$ and $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$, we have $\|\varphi(T)\|\leq 1$, and so $1-\varphi(T)^{*}\varphi(T)\geq 0$.
-In conjunction with the inequality (6), this implies that $(1-\varphi(T)^{*}\varphi(T))x=0$, and hence $\|\varphi(T)x\|^{2}=\|x\|^{2}$.
-Since $x\neq 0$, $x$ is a maximizing vector for $\varphi(T)$ and $\|\varphi(T)\|=1$.
+In conjunction with the equality (6), this implies that $(1-\varphi(T)^{*}\varphi(T))x=0$, and hence $\|\varphi(T)x\|^{2}=\|x\|^{2}$.
+Since $\neq 0, xx$ is a maximizing vector for $\varphi(T)$ and $\|\varphi(T)\|=1$.
 By Theorem 5.12, for the rational function $\varphi$ there exists a model $(\mathcal{M},u),$ where $u:R_{\delta}\to\mathcal{M}$ is holomorphic, so that
 $1-\overline{\varphi(\mu)}\varphi(\lambda)=\Big\langle\,\big(1-E(\mu)^{*}E(\lambda)\big)\ u(\lambda)\,,u(\mu)\,\Big\rangle_{\mathcal{M}}\ \text{for}\ \lambda,\mu\in R_{\delta},$ (6.19)
-where $\dim\mathcal{M}\leq 2n$. Since $T\in\mathcal{F}_{\rm dp}(\delta,\lambda)$, $T$ satisfies
+where $\dim\mathcal{M}\leq 2n$. Since $\in\mathcal{F}_{\rm dp}(\delta,\lambda), TT$ satisfies
 $\sigma(T)=\{\lambda_{1},\ldots,\lambda_{n}\}\subset R_{\delta}.$
 Thus, by the Riesz-Dunford functional calculus, $\varphi(T)$ is well defined and, by the hereditary functional calculus,
-$1-\varphi(T)^{*}\varphi(T)=u(T)^{*}\big{(}1-E(T)^{*}E(T)\big{)}\ u(T).$
+$1-\varphi(T)^{*}\varphi(T)=u(T)^{*}\big(1-E(T)^{*}E(T)\big)\ u(T).\qed$
+7. Declarations
+NSF and EPSRC grants.
+Agler was partially supported by National Science Foundation Grants DMS 1361720 and 1665260.
+Lykova and Young were partially supported by the Engineering and Physical Sciences grant EP/N03242X/1.
+Conflict of interest. The authors have no Conflict of interest to declare that are relevant to the content of this article.
+Data availability statement. No data were collected, generated or consulted in connection with this research.